### v1

# For Better or For Worse? Learning Minimum Variance Features With Label Augmentation

Abstract
Data augmentation has been pivotal in successfully training deep learning models on classification tasks over the past decade. An important subclass of data augmentation techniques - which includes both label smoothing and Mixup - involves modifying not only the input data but also the input label during model training. In this work, we analyze the role played by the label augmentation aspect of such methods. We prove that linear models on linearly separable data trained with label augmentation learn only the minimum variance features in the data, while standard training (which includes weight decay) can learn higher variance features. An important consequence of our results is negative: label smoothing and Mixup can be less robust to adversarial perturbations of the training data when compared to standard training. We verify that our theory reflects practice via a range of experiments on synthetic data and image classification benchmarks.
1 Introduction
The training and fine-tuning procedures for current state-of-the-art (SOTA) computer vision models rely on a number of different data augmentation schemes applied in tandem (Yu et al., 2022; Wortsman et al., 2022; Dehghani et al., 2023). While some of these methods involve only transformations to the input training data - such as random crops and rotations (Cubuk et al., 2019) - a non-trivial subset of them also apply transformations to the input training label.
Perhaps the two most widely applied data augmentation methods in this subcategory are label smoothing (Szegedy et al., 2015) and Mixup (Zhang et al., 2018). Label smoothing replaces the one-hot encoded labels in the training data with smoothed out labels that assign non-zero probability to every possible class (see Section 2 for a formal definition). Mixup similarly smooths out the training labels, but does so via introducing random convex combinations of data points and their labels. As a result, Mixup modifies not only the training labels but also the training inputs. The general principles of label smoothing and Mixup have been extended to create several variants, such as Structural Label Smoothing (Li et al., 2020), Adaptive Label Smoothing (Wang et al., 2021), Manifold Mixup (Verma et al., 2019), CutMix (Yun et al., 2019), PuzzleMix (Kim et al., 2020), SaliencyMix (Uddin et al., 2020), AutoMix (Liu et al., 2021), and Noisy Feature Mixup (Lim et al., 2022).
Due to the success of label smoothing and Mixup-based approaches, an important question on the theoretical side has been understanding when and why these data augmentations improve model performance. Towards that end, several recent works have studied this problem from the perspectives of regularization (Guo et al., 2019; Carratino et al., 2020; Lukasik et al., 2020; Chidambaram et al., 2021), adversarial robustness (Zhang et al., 2020), calibration (Zhang et al., 2021; Chidambaram & Ge, 2023), feature learning (Chidambaram et al., 2023; Zou et al., 2023), and sample complexity (Oh & Yun, 2023).
Although the connection between the label smoothing and Mixup losses has been noted in both theory and practice (Carratino et al., 2020), there has not been (to the best of our knowledge) a unifying theoretical perspective on why the models trained using these losses exhibit similar behavior. In this paper, we aim to provide such a perspective by extending the ideas from previous feature learning analyses to the aspect shared in common between both Mixup and label smoothing: label augmentation. Our analysis allows us to identify a clear failure mode shared by both Mixup and label smoothing - that they can ignore higher variance features - which we also verify empirically.
1.1 Outline of Main Contributions
Our main contribution can be summarized as,
For linearly separable data distributions that can be learned using any one of multiple features, models trained with Mixup or label smoothing learn only the minimum variance feature.
In order to work up to this claim, we provide the necessary background on Mixup and label smoothing and formalize the notion of data distributions with multiple separating features in Section 2. Our definition draws heavily from the idea of multi-view data introduced by Allen-Zhu & Li (2021), but is considerably simpler as it does not involve patches or feature noise.
We then prove our main results in Section 3, where we show that linear models trained with Mixup or label smoothing on the aforementioned type of binary class data distribution are correlated only with the minimum variance feature in the data. On the other hand, we also show that linear models trained using just empirical risk minimization (ERM) and weight decay (so as to ensure a unique solution) have non-trivial correlation with the high variance feature in the data.
A key aspect of our analysis is that we can show a clear separation in terms of feature learning between Mixup/label smoothing and ERM in the setting of just linear models and linearly separable data distributions, whereas previous feature learning analyses of Mixup/similar (Chidambaram et al., 2023; Zou et al., 2023) have relied on data distributions that are explicitly constructed to not be learnable by linear models and have an asymptotically large number of classes. As a result, our analysis is both simpler than existing analyses and shows that nonlinear training dynamics are not necessary to separate Mixup/label smoothing and ERM with respect to feature learning.
We also verify that our theoretical results extend to practice across several experimental settings. In Section 4.1, we show that for synthetic data following our theoretical assumptions, logistic regression trained with Mixup or label smoothing learns a weight vector that is effectively uncorrelated with the high variance feature in the data. We also show that this behavior extends beyond synthetic data in Section 4.2, where we train logistic regression models on binary classification versions of CIFAR-10 and CIFAR-100 modified to have adversarial pixels in the training data that completely identify each class. In this case, the Mixup and label smoothing logistic regression models do not generalize at all due to having learned only the adversarial pixels in the training data.
Finally, we show in Section 4.3 that these observations still apply even when we consider ResNets trained on the full versions of CIFAR-10 and CIFAR-100; for these experiments we follow the experimental setup introduced by Chidambaram et al. (2023) in which the training data is modified to have extra low variance features via additional channels. Our results provide a contrasting perspective to that of Chidambaram et al. (2023) - Mixup (and label smoothing) can end up learning only the low variance features in the training data, whereas ERM can avoid this problem.
1.2 Related Work
Label Smoothing. Since being introduced by Szegedy et al. (2015), label smoothing continues to be used to train SOTA vision (Wortsman et al., 2022; Liu et al., 2022), translation (Vaswani et al., 2017; Team et al., 2022), and multi-modal models (Yu et al., 2022). Attempts to understand when and why label smoothing is effective can be traced back to the influential empirical analyses of Pereyra et al. (2017) and Müller et al. (2020), which respectively relate label smoothing to entropy regularization and show that it can lead to more closely clustered learned representations, with the latter work further showing that these clustered representations can be responsible for decreased distillation performance.
On the theoretical side, Lukasik et al. (2020) study the relationship between label smoothing and loss correction techniques used to handle label noise, and show that label smoothing can be effective for mitigating label noise. Liu (2021) extends this line of work by analyzing how label smoothing can outperform loss correction in the context of memorization of noisy labels. Wei et al. (2022) further show that negative label smoothing can outperform traditional label smoothing in the presence of label noise. Xu et al. (2020) provide an alternative theoretical perspective, showing that label smoothing can improve convergence of stochastic gradient descent. However, to the best of our knowledge, there are no preexisting theoretical works attempting to understand how label smoothing affects feature learning.
Mixup. Similar to label smoothing, Mixup (Zhang et al., 2018) and its aforementioned variants have also played an important role in the training of SOTA vision (Wortsman et al., 2022; Liu et al., 2022), text classification (Sun et al., 2020), translation (Li et al., 2021), and multi-modal models (Hao et al., 2023), often being applied alongside label smoothing. Initial work on understanding Mixup studied it from the perspective of introducing a data-dependent regularization term to the empirical risk (Carratino et al., 2020; Zhang et al., 2020; Park et al., 2022), with Zhang et al. (2020) and Park et al. (2022) showing that this regularization effect can lead to improved Rademacher complexity.
On the other hand, Guo et al. (2019) and Chidambaram et al. (2021) show that the regularization terms introduced by the Mixup loss can also have a negative impact due to the fact that Mixup-augmented points may coincide with existing training data points, leading to models that fail to minimize the original risk. Mixup has also been studied from the perspective of calibration (Thulasidasan et al., 2019), with theoretical results (Zhang et al., 2021; Chidambaram & Ge, 2023) showing that Mixup training can prevent miscalibration issues arising from ERM (even when post-training calibration methods are used with ERM).
Very recently, Oh & Yun (2023) studied Mixup in a similar context to our work (binary classification with linear models), and showed that Mixup can significantly improve the sample complexity required to learn the optimal classifier when compared to ERM. Also recently, and most closely related with our results in this paper, Mixup has been studied from a feature learning perspective, with Chidambaram et al. (2023) and Zou et al. (2023) both showing that Mixup training can learn multiple features in a generative, multi-view (Allen-Zhu & Li, 2021) data model despite ERM failing to do so. We provide a more detailed comparison to these works when we discuss our main results in Section 3.
General Data Augmentation. General data augmentation techniques have been a mainstay in image classification since the rise of deep learning models for such tasks (Krizhevsky et al., 2012). As a result, there is an ever-growing body of theory (Bishop, 1995; Dao et al., 2019; Wu et al., 2020; Hanin & Sun, 2021; Rajput et al., 2019; Yang et al., 2022; Wang et al., 2022; Chen et al., 2020; Mei et al., 2021) aimed at addressing broad classes of augmentations such as those resulting from group actions (i.e. rotations and reflections). Recently, Shen et al. (2022) also studied a class of linear data augmentations from a feature learning perspective, once again using a data model based on the multi-view data of Allen-Zhu & Li (2021). This work is in the same vein as that of Zou et al. (2023) and Chidambaram & Ge (2023), although the augmentation considered is not comparable to Mixup.
2 Preliminaries
Notation. Given $n\in\mathbb{N}$, we use $[n]$ to denote the set $\{1,2,...,n\}$. For a vector $x\in\mathbb{R}^{d}$ and a subset $S\subset[d]$, we use $x_{S}\in\mathbb{R}^{\absolutevalue{S}}$ to denote the restriction of $x$ to only those indices in $S$, and also use $x_{i}$ to denote the $i^{\text{th}}$ coordinate of $x$. The same notation also applies to matrices; i.e. for a matrix $T$ we use $T_{S}$ to denote the restriction of the rows of $T$ to only those dimensions in $S$. Additionally, we use $\succ$ to denote the partial order over positive definite matrices and $\mathrm{Id}$ to denote the identity matrix. We use $\norm{\cdot}$ to indicate the Euclidean norm on $\mathbb{R}^{d}$. For two sets of points $A,B\subset\mathbb{R}^{d}$ we use $d(A,B):=\min_{x\in A,y\in B}\|x-y\|$ to denote the distance between $A$ and $B$. For a function $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ we use $g^{i}(x)$ to denote the $i^{\text{th}}$ coordinate function of $g$. For a probability distribution $\pi$ we use $\mathrm{supp}(\pi)$ to denote its support. Additionally, if $\pi$ corresponds to the joint distribution of two random variables $X$ and $Y$ (i.e. data and label), we use $\pi_{X}$ and $\pi_{Y}$ to denote the respective marginals, and $\pi_{X\mid Y=y}$ and $\pi_{Y\mid X=x}$ to denote the conditional distributions.
Our theory will focus on binary classification problems in $\mathbb{R}^{d}$ in which a subset of the input dimensions correspond to a low variance feature, and the complementary dimensions correspond to a higher variance but more separable feature. This setup is intended to mimic standard explorations of robustness of image classification models. For example, if each class in the data has a small set of constant, identifying pixels in every training point (either introduced adversarially or as an artifact of the domain being considered), does a model trained on this data learn to predict using only the identifying pixels or the whole image?
The goal in such cases is typically to also learn the features present in the non-identifying pixels (i.e. the higher variance features). We will show, however, that doing any kind of label augmentation (label smoothing or Mixup) in this low variance/high variance feature setup will lead to learning a model that has only learned the low variance feature, while standard training will lead to having non-trivial correlation with the high variance feature.
It is also natural to ask whether this phenomenon is purely a result of the fact that augmented labels (i.e. not 0 or 1) prevent the learned model weights from going to infinity, which has been pointed out to be crucial in previous existing analyses of Mixup (Chidambaram et al., 2023). We answer this question in the negative by also showing that training using ERM with weight decay still learns the high variance feature in the data, so the variance minimization effect is more than just a weight norm constraint.
2.1 Data Model
We begin with the generative model of the data that we will work with. For simplicity, we consider the low variance feature to have zero variance conditional on the label $y$ it is not difficult to extend our results to the case where the variance of this feature is simply much less than that of the high variance feature, but considering zero conditional variance yields cleaner proofs. We also consider balanced classes to further simplify the proofs, but our results will hold so long as each class appears with some non-zero probability.
Definition 2.1.
[Data Distribution] Let $\mathcal{L}\subset[d]$ and $\mathcal{H}=\mathcal{L}^{c}$ be the sets denoting the low variance feature and high variance feature in the data respectively. We define $\pi$ to be any distribution on $\mathbb{R}^{d}\times\{-1,1\}$ with $E_{x\sim\pi_{X}}[x]=0$ and $\pi_{Y}(y)=1/2$ for both $y\in\{-1,1\}$, along with the following properties:
  1. 1.
    (Zero Conditional Variance in $\mathcal{L}$) For $\pi_{X}$-a.e. $x\in\mathrm{supp}(\pi_{X\mid Y=y})$, we have that there exists a $v_{y}^{*}\in\mathbb{R}^{\absolutevalue{\mathcal{L}}}$ such that $x_{\mathcal{L}}=v_{y}^{*}$ with $\norm{v_{1}^{*}-v_{-1}^{*}}>0$.
  2. 2.
    (Non-zero Conditional Variance in $\mathcal{H}$) Let $\mu_{y}=\mathbb{E}[X\mid Y=y]$ and $\Sigma_{y}=\mathbb{E}[(x-\mu_{y})(x-\mu_{y})^{\top}\mid Y=y]$, then $\Sigma_{y,\mathcal{H}}$ (the restriction of $\Sigma_{y}$ to $\mathcal{H}$) is positive definite.
  3. 3.
    (More Separable High Variance Feature) Let $A_{y}=\{x_{\mathcal{H}}:x\in\mathrm{supp}(\pi_{X\mid Y=y})\}$. Then $A_{1},A_{-1}$ are compact, convex, and disjoint and satisfy $d(A_{1},A_{-1})>\norm{v_{1}^{*}-v_{-1}^{*}}$.
Since Definition 2.1 corresponds to a linearly separable data distribution, we can restrict our attention to linear models of the form $g_{w}(x)=\sigma(w^{\top}x)$, where $\sigma(\cdot)$ is the sigmoid function. We can then define the binary cross-entropy with weight decay as:
$\ell_{\lambda}(w)$ $=\mathbb{E}_{(x,y)\sim\pi}\left[-\log g_{w}(yx)\right]+\frac{\lambda}{2}\norm{w}^{2}$ (2.1)
With $\ell_{0}(w)$ thereby indicating the regular binary cross-entropy without weight decay. In what follows, we will work with the population form of the loss as defined in Equation (2.1), since for $\lambda>0$ we have that $\ell_{\lambda}(w)$ is strictly convex and therefore we may simply analyze its unique minimizer as opposed to worrying about training dynamics.
Although $\ell_{0}(w)$ will not have a unique minimizer, our results will still apply to the common case where $\ell_{0}(w)$ is minimized by scaling the max-margin solution, which we define below.
Definition 2.2.
[Max-Margin Solution] The max-margin solution $w^{*}$ with respect to $\pi$ is defined to be:
$w^{*}$ $=\operatorname*{argmin}_{w\in\mathbb{R}^{d}}\norm{w}^{2}$
[…]
2.2 Label Smoothing
Label smoothing as introduced by Szegedy et al. (2015) is intended to work with multi-class classification problems in which the data distribution $\pi$ is typically defined on $\mathbb{R}^{d}\times[k]$. In this case, we work with models $g:\mathbb{R}^{d}\to\Delta^{k}$ (the probability simplex in $\mathbb{R}^{k}$), and the cross-entropy loss is simply $\ell(g)=\mathbb{E}_{(x,y)\sim\pi}[-\log g^{y}(x)]$.
The label-smoothed cross-entropy is then obtained by replacing the one-hot encoded label $y$ with a mixture of the original label and the uniform distribution over $[k]$. Namely, the label-smoothed loss with mixing hyperparameter $\alpha\in[0,1]$ is defined to be:
$\ell_{\mathrm{LS},\alpha}(g)=-\mathbb{E}_{(x,y)\sim\pi}\bigg{[}$ $(1-\alpha)\log g^{y}(x)+\frac{\alpha}{k}\sum_{i=1}^{k}\log g^{i}(x)\bigg{]}$ (2.3)
For the purposes of our theory, however, we will be working with models of the form $g_{w}(x)=\sigma(w^{\top}x)$ and labels $y\in\{-1,1\}$, so we will consider the binary label-smoothed loss defined as:
$\ell_{\mathrm{LS},\alpha}(w)=-\mathbb{E}_{(x,y)\sim\pi}\bigg{[}$ $\left(1-\frac{\alpha}{2}\right)\log g_{w}(yx)+\frac{\alpha}{2}\log g_{w}(-yx)\bigg{]}$ (2.4)
Note that similar to Equation (2.1) the latter is defined as a function of the weight vector $w$ instead of the full model $g$.
2.3 Mixup
Similar to label smoothing, Mixup (Zhang et al., 2018) was also introduced in the context of multi-class classification. In addition to augmenting the one-hot encoded labels $y$, Mixup also augments the input data $x$ by combining both with another data point $(x^{\prime},y^{\prime})$ via a random convex combination. Both of these transformations are done according to a mixing distribution $\mathcal{D}_{\lambda}$ (which is a hyperparameter) whose support is contained in $[0,1]$. To simplify notation, we will use $x_{1:n}$ and $y_{1:n}$ to denote multiple data points $x_{1},...,x_{n}$ and their corresponding labels $y_{1},...,y_{n}$. Additionally, in the multi-class setting with $g:\mathbb{R}^{d}\to\Delta^{k}$, we also define:
$h_{\lambda}(g,x_{1:2},y_{1:2})$ $=\lambda\log g^{y_{1}}(z_{\lambda})+(1-\lambda)\log g^{y_{2}}(z_{\lambda})$ (2.5)
$\text{where}\quad z_{\lambda}$ $=\lambda x_{1}+(1-\lambda)x_{2}$ (2.6)
We may then define the Mixup cross-entropy:
$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(g)=\mathbb{E}_{(x_{1:2},y_{1:2})\sim\pi,\lambda\sim\mathcal{D}_{\lambda}}\left[-h_{\lambda}(g,x_{1:2},y_{1:2})\right]$ (2.7)
Where $(x_{1:2},y_{1:2})\sim\pi$ is shorthand for $(x_{1},y_{1}),(x_{2},y_{2})\sim\pi$. As before, since our theory focuses on binary classification and linear models $g_{w}$, we may identically define the binary version $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ of Equation (2.7) except for redefining $h_{\lambda}$ to be:
$h_{\lambda}(w,x_{1:2},y_{1:2})$ $=\lambda\log g_{w}(y_{1}z_{\lambda})+(1-\lambda)\log g_{w}(y_{2}z_{\lambda})$ (2.8)
Again, observe that we have redefined $h_{\lambda}$ to directly be a function of $w$ instead of $g$.
3 Main Theoretical Results
Having established the necessary background, we now state our main results. Full proofs of all results can be found in Appendix A. We first show that the unique minimizer $w^{*}$ of the standard risk with weight decay (as defined in Equation (2.1)) has non-trivial correlation with the high variance dimensions of the input $x$.
Theorem 3.1.
Let $w^{*}$ be the unique minimizer of $\ell_{\lambda}(w)$ for $\lambda>0$. Then $\norm{w^{*}_{\mathcal{H}}}^{2}\geq\frac{1}{2}\norm{w^{*}}^{2}$. In other words, the optimal solution under weight decay has significant weight in the high variance dimensions $\mathcal{H}$.
Proof Sketch. We can orthogonally decompose the optimal solution $w^{*}$ in terms of the unit normal max-margin directions $u_{1}^{*}$ and $u_{2}^{*}$ separating $x_{\mathcal{L}}$ and $x_{\mathcal{H}}$ respectively. Then, using the fact that $y\left\langle u_{2}^{*},x_{\mathcal{H}}\right\rangle>y\left\langle u_{1}^{*},x_{\mathcal{L}}\right\rangle$ for all $(x,y)\in\mathrm{supp}(\pi)$ (from assumption three in Definition 2.1), we can claim that $w^{*}$ must have greater weight associated with $u_{2}^{*}$ than $u_{1}^{*}$, as otherwise we can decrease $\ell_{\lambda}(w^{*})$ by moving some weight from $u_{1}^{*}$ to $u_{2}^{*}$.
We cannot immediately extend the result of Theorem 3.1 to the binary cross-entropy $\ell_{0}(w)$ without weight decay, since there is no unique minimizer of $\ell_{0}$. However, for linear models trained with gradient descent (as is often done in practice), it is well-known that the learned model converges in direction to the max-margin solution (Soudry et al., 2018; Ji & Telgarsky, 2020). For this case, the proof technique of Theorem 3.1 readily extends and we obtain the following corollary.
Corollary 3.2.
If $w^{*}$ is the max-margin solution to $\ell_{0}(w)$, then the result of Theorem 3.1 still holds.
On the other hand, as alluded to earlier, this phenomenon does not occur when considering the label smoothing and Mixup losses $\ell_{\mathrm{LS},\alpha}(w)$ and $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$. Indeed, we can show:
Theorem 3.3.
Every minimizer $w^{*}$ of $\ell_{\mathrm{LS},\alpha}(w)$ for $\alpha>0$ satisfies $\norm{w^{*}_{\mathcal{H}}}=0$, so the optimal solutions under label smoothing have no weight in the high variance dimensions.
Proof Sketch. The optimal prediction at a label-smoothed point is either $21-\alpha/$ 2 or $\alpha/2$ (depending on whether the label is $1$ or $-1$ respectively). We can thus construct a predictor that is optimal at every point by choosing $w$ to have the appropriate correlation with $v_{1}^{*}$ and having no correlation with $x_{\mathcal{H}}$. That there are no equally good predictors that have correlation with $x_{\mathcal{H}}$ can be shown by appealing to convexity.
Analogously, we have the following result for Mixup:
Theorem 3.4.
Every minimizer $w^{*}$ of $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ satisfies $\norm{w^{*}_{\mathcal{H}}}=0$ for any distribution $\mathcal{D}_{\lambda}$ with $\mathrm{supp}(\mathcal{D}_{\lambda})\subseteq[0,1]$ that is not a point mass on either 0 or 1.
Proof Sketch. The proof setup is similar to that of Theorem 3.3, i.e. we try to find a stationary point of $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$ that is zero in the high variance dimensions. The difficulty in the Mixup analysis is that it is not possible to directly construct a point-wise optimal predictor like in the proof of Theorem 3.3 due to the different possibilities for mixed points (and mixed labels), and consequently there is no obvious construction for a stationary point. However, we can analyze the limiting behavior of the Mixup gradient with respect to the weight correlation with the low variance dimensions and still show that such a stationary point must exist.
Comparisons to existing results. We are not aware of any existing results on feature learning for label smoothing; prior theoretical work largely focuses on the relationship between label smoothing and learning under label noise, which is orthogonal to the perspective we take in this paper. Although feature learning results exist for Mixup (Chidambaram et al., 2023; Zou et al., 2023), the pre-existing results are constrained to the case of mixing using point mass distributions and only prove a separation between Mixup and unregularized ERM, whereas our results work for arbitrary mixing distributions (that don’t coincide with ERM) and also separate Mixup from ERM with weight decay. This greater generality comes at the cost of considering only linear models and linearly separable data, however, as both Chidambaram et al. (2023) and Zou et al. (2023) consider non-linearly separable data and the training dynamics of 2-layer neural networks.
Additionally, our results can be viewed as generalizing the observations made by Chidambaram et al. (2023), which were that Mixup can learn multiple features in the data when doing so decreases the variance of the learned predictor. In our case, we directly show that Mixup will learn the low variance feature in the data as opposed to the high variance feature. Our results also do not contradict the observations of Zou et al. (2023), which were that Mixup can learn both a “common” feature and a “rare” feature in the data; in their setup the common/rare features are fixed (zero variance) per class and concatenated with high variance noise.
4 Experiments
(a) Weight Decay
(b) Label Smoothing
(c) Mixup
Figure 1: Final weight norm of logistic regression model in the high variance dimensions ($\norm{w_{\mathcal{H}}}$) for various hyperparameter settings on the synthetic data distribution of Section 4.1.
We now verify the predictions of our theoretical results in Section 3 across a variety of experimental setups. All experiments in this section were conducted on a single P100 GPU using PyTorch (Paszke et al., 2019) for model implementation. All reported results correspond to means over 5 training runs, and shaded regions in the figures correspond to 1 standard deviation bounds.
4.1 Synthetic Data
We first consider the case of training logistic regression models on a synthetic data distribution that exactly follows the assumptions of Definition 2.1.
Definition 4.1 (Synthetic Data).
We define a distribution $\pi_{\gamma}$ parameterized by $\gamma$ (where $0<\gamma\ll 1$) on $\mathbb{R}^{d}\times\{-1,1\}$ with $\pi_{\gamma,Y}(1)=1/2$ and $x\sim\pi_{\gamma,X\mid Y=y}$ satisfying:
  1. 1.
    First $\lfloor d/2\rfloor$ dimensions are constant but small. We have $x_{i}=\gamma y$ for $i\leq\lfloor d/2\rfloor$.
  2. 2.
    Last $d-\lfloor d/2\rfloor$ dimensions are high variance. We have $x_{i}\sim\mathrm{Uniform}([y,100y])$ for $i>\lfloor d/2\rfloor$.
In other words, we consider data where (conditional on the label) the first half of the dimensions are fixed (corresponding to $\mathcal{L}$) and the second half are i.i.d. high variance uniform (corresponding to $\mathcal{H}$). We sample $n=5000$ data points according to Definition 4.1 with $d=10$ and $\gamma=0.1$ and train logistic regression models across a range of settings for weight decay, label smoothing, and Mixup. [^1] The choice of $\gamma$ here and throughout the rest of this section is relatively arbitrary; we verified that our experimental phenomena are unchanged for other scales such as $10^{-2},10^{-3},\text{ and }10^{-4}$.
Namely, we consider 20 uniformly spaced values in $[0,0.1]$ for the weight decay $\lambda$ parameter and in $[0,0.75]$ for the label smoothing $\alpha$ parameter, where the upper bound for the label smoothing parameter space is obtained from the experiments of Müller et al. (2020). For Mixup, we fix the mixing distribution to be the canonical choice of $\mathrm{Beta}(\alpha,\alpha)$ introduced by Zhang et al. (2018) and consider 20 uniformly spaced $\alpha$ values in $[0,8]$ (with $\alpha=0$ corresponding to ERM), which effectively covers the range of Mixup hyperparameter values used in practice.
We train all models for 100 epochs using Adam with the standard hyperparameters of $\beta_{1}=0.9,\beta_{2}=0.999$, a learning rate of $5*10^{-3}$, and a batch size of 500. At the end of training, we compute $\norm{w_{\mathcal{H}}}$ (i.e. the norm of the weight vector in the last 5 dimensions) for each trained model. For each model setting, we report the mean and standard deviation of $\norm{w_{\mathcal{H}}}$ over 5 training runs in Figure 1.
As can be seen from the results, the weight decay models always have non-trivial values for $\norm{w_{\mathcal{H}}}$ (even for large values of $\lambda$), whereas the label smoothing and Mixup models very quickly converge to a $\norm{w_{\mathcal{H}}}$ of effectively zero as their respective hyperparameters move away from the ERM regime ($\alpha=0$ in both cases). This matches the behavior predicted by Theorems 3.1 to 3.4.
It is natural to ask to what extent these results depend on the fact that the low variance features in Definition 4.1 (the first half of the dimensions) actually have zero variance conditional on the label. As we mentioned in our theoretical results, it is sufficient to assume that the variance of the dimensions in $\mathcal{L}$ is simply much less than that of $\mathcal{H}$, and the fact that we consider zero variance is merely for simplifying purposes. We verify that this is true in practice by re-running the above experiments but modifying the dimensions in $\mathcal{L}$ to have non-zero variance by adding uniformly distributed noise on the scale of $\gamma/10$. The results remain unchanged; the corresponding plots can be found in Section B of the Appendix.
4.2 Binary Image Classification
(a) Weight Decay
(b) Label Smoothing
(c) Mixup
Figure 2: Logistic regression final test errors for various hyperparameter settings on the binary classification versions of CIFAR-10 and CIFAR-100 from Section 4.2.
Although the experiments of Section 4.1 constitute a direct verification of our theory, in practice the low and high variance features are typically not as obviously partitioned as in Definition 4.1 and the metric we care about is usually generalization error as opposed to the strength of correlation with some part of the data. As such, we now relax our strict adherence to the assumptions of Definition 2.1 and consider binary classification on image data.
To do so, we reduce the standard benchmarks of CIFAR-10 and CIFAR-100 (Krizhevsky, 2009) to binary classification tasks by maintaining only the first two classes (“airplane” and “automobile” for CIFAR-10, “apple” and “aquarium fish” for CIFAR-100) and assigning them the labels of -1 and 1 respectively. This leads to 10000 training examples and 2000 test examples in the case of CIFAR-10, and 1000 training examples and 200 test examples in the case of CIFAR-100.
We preprocess the training data to have mean zero and variance one along every image channel (and apply the same transform to the test data), and then modify the training data such that the top left value (corresponding to the “R” channel of the top left pixel in the original image) in the tensor representation of the training input is replaced by $y\gamma$, where $\gamma=0.1$ as in Section 4.1. This ensures that the training data is separable in the first dimension of the data, although unlike in the case of Definition 4.1 we make no modifications to the other dimensions of the data to ensure separability outside the first dimension. We leave the test data unchanged - our goal is to determine whether models trained on the modified training data can learn more than just the single identifying dimension.
We follow the exact same logistic regression training setup as in Section 4.1, and sweep over the same hyperparameter values for weight decay, label smoothing, and Mixup. We report the mean test error along with one standard deviation error bounds over five training runs for each model setting in Figure 2.
The results mirror our observations in Section 4.1: both the label smoothing and Mixup models have high test error for all settings (in this case, ERM performs poorly as well) while weight decay achieves a significantly lower test error for all $\lambda>0$. The higher variance in the reported CIFAR-100 test errors can almost certainly be attributed to the fact that our binary version of CIFAR-100 consists of only 1000 training points and 200 points.
(a) CIFAR-10
(b) CIFAR-100
Figure 3: Comparison of norm ratio between first dimension (synthetically modified in the training data) and remaining dimensions (left unchanged) for trained logistic regression weight vector for the 20 different hyperparameter settings of weight decay, label smoothing, and Mixup.
Our results correspond once again to label smoothing and Mixup learning to use only the zero variance, identifying dimension of the input, as we expect from our theory. Indeed, we verify this in Figure 3 where we plot the ratio $\norm{w_{1}}/\norm{w_{[d]\setminus\{1\}}}$ (i.e. the ratio between the norm of the trained model weight vector in the first dimension and the remaining dimensions). We observe that for both CIFAR-10 and CIFAR-100, across all the non-ERM hyperparameter settings (refer to Section 4.1 for the range of hyperparameters) that both label smoothing and Mixup place much greater weight in the first dimension than weight decay.
As with the experiments of Section 4.1, we again check that the results here do not rely on the fact that the identifying dimension introduced into the training data has zero variance conditional on the label. As before, introducing uniformly distributed noise on the scale of $\gamma/10$ to the identifying dimension does not change our results, and the corresponding plots can again be found in Section B of the Appendix.
4.3 Multi-Class Image Classification
(a) Weight Decay ($\lambda=0.01$)
(b) Label Smoothing ($\alpha=0.1$)
(c) Mixup ($\alpha=1$)
Figure 4: ResNet-18 test error at the end of each epoch of training on the modified versions of CIFAR-10 and CIFAR-100 from Section 4.3, using fixed settings for weight decay, label smoothing, and Mixup.
Thus far we have considered training logistic regression models to stay within the confines of our theory, but we now investigate whether our observations extend to the more practical setting of training nonlinear models (i.e. deep neural networks) on multi-class image classification data.
As in Section 4.2, we consider CIFAR-10 and CIFAR-100, but this time maintain all of the original classes in both datasets. We once again preprocess the training data to have mean zero and variance one, and apply this transform to the test data as well. Unfortunately, unlike in Section 4.2, introducing a zero variance, identifying feature into each class in the training data would now require modifying a more significant portion of each image (for example, a $10\times 10$ patch for CIFAR-100 to identify each class).
Figure 5: Visualization of data modification applied to CIFAR-10/CIFAR-100 in Section 4.3.
So as to preserve most of the original image content, as well as to have a more direct comparison to prior work, we choose to adopt a similar experimental setup to that of Chidambaram et al. (2023). Namely, we assign a random vector sampled from $\mathcal{N}(0,\gamma^{2}\mathrm{Id}_{d})$ (where $d$ is the dimension of the data) to each class and concatenate this class vector along the channels dimension to every training data point for that class. A visualization of this transformation is shown in Figure 5. For the test data points, we concatenate an all-zero vector of the same size so as to preserve the dimensionality across train and test, as was done by Chidambaram et al. (2023). We take $\gamma=0.1/\sqrt{d}$, so as to account for the dimensionality of the concatenated feature.
For both datasets, we train ResNet-18 models (He et al., 2015) for 100 epochs using Adam with the standard hyperparameters of $\beta_{1}=0.9,\beta_{2}=0.999$, a learning rate of $10^{-3}$, and a batch size of 500. Due to the higher compute cost of training these models compared to the logistic regression models of Sections 4.1 and 4.2, we do not perform a full hyperparameter sweep but instead use known well-performing settings for weight decay, label smoothing, and Mixup. Namely, we use $\lambda=0.01$ for weight decay, $\alpha=0.1$ for label smoothing (Müller et al., 2020), and $\alpha=1$ in $\mathrm{Beta}(\alpha,\alpha)$ for Mixup (Zhang et al., 2018). Since we are now dealing with nonconvex training dynamics, we report the mean test error over five runs at the end of each epoch with one standard deviation error bounds for each model setting in Figure 4.
As can be seen from the plots, both label smoothing and Mixup initially decrease the test error in the initial stages of training, but then overfit to the concatenated zero variance features. On the other hand, the test error for weight decay decreases over the course of training until effectively converging. This once again does not contradict what was observed by Chidambaram et al. (2023), as in their case their concatenated features were weighted alongside with the original image data so that variance could be decreased by learning both.
We also considered the case of low variance concatenated features as opposed to zero variance concatenated features in the training data. Similar to the results of Sections 4.1 and 4.2, the label smoothing and Mixup overfitting phenomenons persist even when we introduce variance to the concatenated features. However, we found that unless the introduced variance is much smaller (i.e. multiple orders of magnitude) than the concatenated feature scale $\gamma$, weight decay also overfits to the concatenated features. This brittleness is possibly an artifact of the chosen experimental setup and/or hyperparameters, but due to compute constraints we did not exhaustively explore this behavior and instead leave it as an open direction for follow-up work extending our results to non-linear/non-convex settings.
5 Conclusion
In summary, we have shown that for linearly separable data distributions in which there exist high variance features and much lower variance features, linear models trained with the data augmentation schemes of label smoothing and Mixup fail to learn the low variance features in both practice and theory, while standard training with weight decay succeeds. The main practical takeaway from this is that label smoothing and Mixup do not necessarily always improve the robustness of models, and their failure to do so may be indicative of certain identifying features in the training data (i.e. adversarially introduced patches). A natural follow-up to the results we have presented is an extension to the case of non-linear models and non-linearly-separable data, which the results of Section 4.3 indicate should be possible.
Acknowledgements
Rong Ge and Muthu Chidambaram are supported by NSF Award DMS-2031849 and CCF-1845171 (CAREER). Muthu would like to thank Annie Tang for thoughtful feedback during the early stages of this project.
Appendix A Full Proofs of Main Results
Here we provide full proofs of all of the results contained in Section 3 of the main paper.
A.1 Proof of Theorem 3.1 and Corollary 3.2
See 3.1
Proof.
First, we claim that $w_{\mathcal{L}}^{*}$ is in the direction of $v_{1}^{*}$, which is the normal vector to the max-margin hyperplane separating $v_{1}^{*}$ and $v_{-1}^{*}$ (since $v_{1}^{*}=-v_{-1}^{*}$ due to the fact that the data distribution has mean zero and $\pi_{Y}(1)=\pi_{Y}(-1)$). Indeed, for any other choice of $w\in\mathbb{R}^{d}$ and any $(x,y)\in\mathrm{supp}(\pi)$, we have:
$y\left\langle w_{\mathcal{L}},x_{\mathcal{L}}\right\rangle=y\left\langle w_{\mathcal{L}},v^{*}_{y}\right\rangle=\left\langle w_{\mathcal{L}},v^{*}_{1}\right\rangle$ (A.1)
As a result, if $w_{\mathcal{L}}$ is not in the direction of $v_{1}^{*}$, we can remove the component orthogonal to $v_{1}^{*}$ (thereby decreasing the norm of the solution) while keeping the inner product with $x_{\mathcal{L}}$ the same. Now let $u_{1}^{*}=v_{1}^{*}/\norm{v_{1}^{*}}$ and $u_{2}^{*}\in\mathbb{R}^{\absolutevalue{\mathcal{H}}}$ similarly denote the unit normal vector to the max-margin hyperplane separating $A_{1}$ and $A_{-1}$. Recall that $\left\langle u_{2}^{*},x_{\mathcal{H}}\right\rangle>\left\langle u_{1}^{*},x_{\mathcal{L}}\right\rangle$ by assumption three in Definition 2.1.
Let us identify $u_{1}^{*}$ and $u_{2}^{*}$ with the corresponding vectors in $\mathbb{R}^{d}$ whose $-\mathcal{L}$ and $-\mathcal{H}$ components are zero respectively, so that $u_{1}^{*}$ and $u_{2}^{*}$ are orthonormal in $\mathbb{R}^{d}$. Then we can decompose the unique minimizer of the loss $w^{*}$ as $\alpha u_{1}^{*}+\beta u_{2}^{*}+\kappa u_{3}^{*}$, where $u_{3}^{*}$ is a vector orthonormal to $u_{1}^{*}$ and $u_{2}^{*}$.
Now suppose that $\alpha>\beta$. Then we claim that replacing $\alpha$ and $\beta$ with $\gamma=\sqrt{(\alpha^{2}+\beta^{2})/2}$ yields a solution with lower loss than $w^{*}$.
To see this, first observe that $2\gamma^{2}=\alpha^{2}+\beta^{2}$, so that the norm of the modified solution is the same as $w^{*}$. This implies that the weight decay penalty term in $\ell_{\lambda}$ is unchanged.
Furthermore, we have that $\gamma\in(\beta,\alpha)$, and that $\gamma-\beta>\alpha-\gamma$. This follows from the fact that $\gamma\geq(\alpha+\beta)/2$ by construction. This, combined with the fact that $y\left\langle u_{2}^{*},x_{\mathcal{H}}\right\rangle>y\left\langle u_{1}^{*},x_{\mathcal{L}}\right\rangle$, then implies for all $(x,y)\in\mathrm{supp}(\pi)$ that:
$y\left\langle\gamma u_{1}^{*}+\gamma u_{2}^{*},x\right\rangle>y\left\langle\alpha u_{1}^{*}+\beta u_{2}^{*},x\right\rangle$ (A.2)
Which contradicts the minimality of $w^{*}$. Therefore, we must have $\alpha\leq\beta$, which gives the desired result. ∎
See 3.2
Proof.
The result follows from the same decomposition approach of Theorem 3.2 and the definition of the max-margin solution (Definition 2.2). Namely, we can move weight from the max-margin direction in $\mathcal{L}$ to the max-margin direction in $\mathcal{H}$ to decrease the norm while satisfying the separation constraint. ∎
A.2 Proof of Theorem 3.3
See 3.3
Proof.
The strategy of the proof will be to construct a stationary point for the label smoothed loss that is correlated only with the dimensions in $\mathcal{L}$, and then argue using Jensen’s inequality that any solution that is correlated with the dimensions in $\mathcal{H}$ will have non-zero variance and thereby worse loss than our constructed stationary point. We first compute the gradient $\gradient_{w}\ell_{\mathrm{LS},\alpha}(w)$ and Hessian $\gradient_{w}^{2}\ell_{\mathrm{LS},\alpha}(w)$:
$\gradient_{w}\ell_{\mathrm{LS},\alpha}(w)$ $=\mathbb{E}_{(x,y)\sim\pi}\left[(g_{w}(yx)-1+\alpha/2)yx\right]$ (A.3)
$\gradient_{w}^{2}\ell_{\mathrm{LS},\alpha}(w)$ $=\mathbb{E}_{x\sim\pi_{X}}\left[g_{w}(yx)(1-g_{w}(yx))xx^{\top}\right]$ (A.4)
Equations (A.3) and (A.4) follow from interchanging the gradient operator with the expectations, which is justified in this case via the dominated convergence theorem because $g_{w}(x)$ is continuously differentiable (with bounded derivative) in each component of $w$, and $x$ has bounded support. Clearly the Hessian is positive semidefinite, so $\ell_{\mathrm{LS},\alpha}(w)$ is convex.
It is also straightforward to construct a stationary point $w^{*}$ of $\ell_{\mathrm{LS},\alpha}(w)$ by choosing $w^{*}_{\mathcal{L}}=\beta v_{1}^{*}$ for appropriate $\beta\in\mathbb{R}$ and $w^{*}_{\mathcal{H}}=0$, since this allows for $g_{w^{*}}(v_{1}^{*})=1-\alpha/2$ and $g_{w^{*}}(-v_{1}^{*})=g_{w^{*}}(v_{-1}^{*})=\alpha/2$ (here $v_{1}^{*}=-v_{-1}^{*}$ since $\pi_{X}$ has mean zero). Now we claim that any $w^{\prime}$ with $\norm{w^{\prime}_{\mathcal{H}}}>0$ has strictly greater loss than $w^{*}$.
To show this, we first observe that $\ell_{\mathrm{LS},\alpha}$ is also convex as a function of $w^{\top}x$ (this is easily check from Equations (A.3) and (A.4)). Additionally, we can restrict our attention to $w^{\prime}$ satisfying $\norm{w^{\prime}}\leq M$ for some $M$, since $\ell_{\mathrm{LS},\alpha}(w^{\prime})\to\infty$ as $\norm{w^{\prime}}\to\infty$ for $\alpha>0$. Therefore, we can consider $w^{\top}x$ to be integrable and apply the conditional form of Jensen’s inequality to $\ell_{\mathrm{LS},\alpha}(w)$:
$(\displaystyle\ell_{\mathrm{LS},\alpha}(w)\geq\sum_{y\in\{-1,1\}}-\pi_{Y}(y)\bigg{(}$$\bigg{(}1-\frac{\alpha}{2}\bigg{)}\log\sigma\big{(}\mathbb{E}[yw^{\top}x\mid Y=y]\big{)}$
$+\frac{\alpha}{2}\log\sigma\big{(}-\mathbb{E}[yw^{\top}x\mid Y=y]\big{)}\bigg{)}$ (A.5)
The RHS of Equation (A.5) is minimized by our choice of $w^{*}$, and in fact we have equality between $\ell_{\mathrm{LS},\alpha}(w^{*})$ and the RHS. Furthermore, since we are dealing with a non-affine function of $w^{\top}x$, we can only achieve equality if $w^{\top}x$ is almost surely constant conditioned on $Y=y$. However, by assumption two in Definition 2.1, the conditional variance of $(w_{\mathcal{H}}^{\prime})^{\top}x_{\mathcal{H}}$ is non-zero, and therefore $\ell_{\mathrm{LS},\alpha}(w^{\prime})>\ell_{\mathrm{LS},\alpha}(w^{*})$. ∎
A.3 Proof of Theorem 3.4
See 3.4
Proof.
We largely follow the same structure as the proof of Theorem 3.3. As before, we compute $\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ and $\gradient_{w}^{2}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ (once again by applying dominated convergence):
$\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\mathbb{E}_{(x_{1},y_{1}),(x_{2},y_{2})\sim\pi}\left[\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\left(\lambda(g_{w}(y_{1}z_{\lambda})-1)y_{1}+(1-\lambda)(g_{w}(y_{2}z_{\lambda})-1)y_{2}\right)z_{\lambda}\right]\right]$ (A.6)
$\gradient_{w}^{2}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\mathbb{E}_{(x_{1},y_{1}),(x_{2},y_{2})\sim\pi}\bigg{[}\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\bigg{[}\bigg{(}\lambda(1-g_{w}(y_{1}z_{\lambda}))g_{w}(y_{1}z_{\lambda})$
$\quad\quad\quad\quad\quad\quad+(1-\lambda)(1-g_{w}(y_{2}z_{\lambda}))g_{w}(y_{2}z_{\lambda})\bigg{)}z_{\lambda}z_{\lambda}^{\top}\bigg{]}\bigg{]}$ (A.7)
It follows from Equation (A.7) that the Hessian is positive semidefinite (since $\lambda\in[0,1]$), so $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ is also convex.
Now it also follows from the above calculations that, after conditioning on $\lambda,y_{1},\text{ and }y_{2}$, the function $h_{\lambda}$ from Equation (2.8) (redefined below to indicate that $y_{1},y_{2}$ are also fixed) is convex in $w^{\top}z_{\lambda}$:
$h_{\lambda,y_{1},y_{2}}(w,z_{\lambda})=-\left(\lambda\log g_{w}(y_{1}z_{\lambda})+(1-\lambda)\log g_{w}(y_{2}z_{\lambda})\right)$ (A.8)
For convenience, we will use $z_{\lambda}\mid y_{1},y_{2}$ to indicate the conditional distribution of $\lambda x_{1}+(1-\lambda)x_{2}$ given $\lambda,y_{1},y_{2}$ as in Equation (A.8). For the moment we suppose that $w^{\top}z_{\lambda}$ is integrable. Then by Fubini’s theorem, conditioning, and Jensen’s inequality we obtain:
$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\mathbb{E}_{y_{1},y_{2}\sim\pi_{Y}}\left[\mathbb{E}_{z_{\lambda}\mid y_{1},y_{2}}\left[h_{\lambda,y_{1},y_{2}}(w,z_{\lambda})\right]\right]\right]$
$\geq\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\mathbb{E}_{y_{1},y_{2}\sim\pi_{Y}}\left[-\left(\lambda\log\sigma(y_{1}\mathbb{E}[w^{\top}z_{\lambda}])+(1-\lambda)\log\sigma(y_{2}\mathbb{E}[w^{\top}z_{\lambda}])\right)\right]\right]$ (A.9)
Equality is only obtained in Equation (A.9) if and only if $w^{\top}z_{\lambda}\mid y_{1},y_{2}$ is almost surely constant, which only occurs if $\norm{w_{\mathcal{H}}}=0$ as before. It remains to show in this case that we can obtain a stationary point with $\norm{w_{\mathcal{H}}}=0$, as then that will imply that any choice of $w$ with $\norm{w_{\mathcal{H}}}>0$ will have strictly worse loss.
To do so, let us consider $w$ with $\norm{w_{\mathcal{H}}}=0$ and $\left\langle w,v_{1}^{*}\right\rangle=u=-\left\langle w,v_{-1}^{*}\right\rangle$. Let $\mathbb{P}_{\lambda}$ denote the probability measure associated with $\mathcal{D}_{\lambda}$. We then define the following two functions, which will correspond to the terms in $\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ that result from mixing $y_{1}=y_{2}$ and $y_{1}\neq y_{2}$.
$f_{1}(u)$ $=\sigma(u)-1$ (A.10)
$f_{2}(u,\lambda)$ $=\sigma((2\lambda-1)u)-\lambda$ (A.11)
Now, recalling that $\mu_{y}=\mathbb{E}[X\mid Y=y]$ and that $\mu_{1}=-\mu_{-1}$ since $\pi_{Y}(1)=\pi_{Y}(-1)$ and $\mathbb{E}_{x\sim\pi_{X}}[x]=0$, we get:
$\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\mathbb{E}_{y_{1},y_{2}\sim\pi_{Y}}\left[\mathbb{E}_{z_{\lambda}\mid y_{1},y_{2}}\left[\left(\lambda(g_{w}(y_{1}z_{\lambda})-1)y_{1}+(1-\lambda)(g_{w}(y_{2}z_{\lambda})-1)y_{2}\right)z_{\lambda}\right]\right]\right]$
$=\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\frac{1}{2}\mathbb{E}_{z_{\lambda}\mid y_{1}=1,y_{2}=1}\left[f_{1}(u)z_{\lambda}\right]+\frac{1}{2}\mathbb{E}_{z_{\lambda}\mid y_{1}=1,y_{2}=-1}\left[f_{2}(u,\lambda)z_{\lambda}\right]\right]$
[…]
Where in the penultimate step we used the fact that the terms that result from conditioning with $y_{1}=y_{2}$ are the same after substituting $\mu_{-1}=-\mu_{1}$, and likewise for the mixed term. Now we consider the limiting behavior of Equation (A.12) as $u\to\pm\infty$. We first focus on the integral term in Equation (A.12), which we can analyze via dominated convergence:
[…]
$\quad+\int_{0.5}^{1}\big{(}\sigma((2\lambda-1)u)-\lambda\big{)}(2\lambda-1)\ d\mathbb{P}_{\lambda}(\lambda)\bigg{)}$
[…]
$\geq 0$ (A.13)
And similarly, the limit as $u\to-\infty$ is bounded above by 0. Thus, from Equation (A.12) we get that:
$\lim_{u\to\infty}\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\beta_{1}\mu_{1}\quad\text{for some}\quad\beta_{1}\geq 0$ (A.14)
$\lim_{u\to-\infty}\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\beta_{2}\mu_{1}\quad\text{for some}\quad\beta_{2}\leq 0$ (A.15)
Since Equation (A.12) is continuous in $u$, it then follows that there exists $u$ such that $\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)=0$, which implies that we can make our choice of $w$ a stationary point. However, there is one subtlety here - the fact that any other $w^{\prime}$ with $\norm{w^{\prime}_{\mathcal{H}}}>0$ has worse loss hinges on the use of Jensen’s inequality above, which requires that we constrain ourselves to the case where $w^{\top}z_{\lambda}$ is integrable.
[…]
For this case, if $\norm{w^{\prime}_{\mathcal{H}}}>0$, then $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w^{\prime})\to\infty$ as $\norm{w^{\prime}}\to\infty$, since we incur infinite loss on the mixed terms $y_{1}\neq y_{2}$. As a result, we can conclude that every minimizer of $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$ (for $\mathcal{D}_{\lambda}$ satisfying the conditions of the theorem) has zero norm in the high variance dimensions, including those minimizers that occur as the norm of the weights tend to infinity. ∎
Appendix B Low Variance Experiments
(a) Weight Decay
(b) Label Smoothing
(c) Mixup
Figure 6: Final weight norm of logistic regression model in the high variance dimensions ($\norm{w_{\mathcal{H}}}$) for various hyperparameter settings on the synthetic data distribution of Section 4.1 modified so that the dimensions in $\mathcal{L}$ have non-zero variance.
(a) Weight Decay
(b) Label Smoothing
(c) Mixup
Figure 7: Logistic regression final test errors for various hyperparameter settings on the binary classification versions of CIFAR-10 and CIFAR-100 from Section 4.2, but with the training data modified to have non-zero variance in the identifying dimension.
Figures 6 and 7 correspond to non-zero variance versions of the experiments of Sections 4.1 and 4.2. In Figure 6, we consider synthetic data as per Definition 4.1 except for that $x_{i}=y(\gamma+\xi)$ where $\xi\sim\mathrm{Uniform}([0,\gamma/10])$ for $i\leq\lfloor d/2\rfloor$ (i.e. the $\mathcal{L}$ dimensions now have non-zero variance). Similarly, Figure 7 considers the same setting as Section 4.2 but with the first dimension of the input replaced with $y(\gamma+\xi)$ instead of $y\gamma$. As can be seen from the results, the introduction of a small amount of variance to these settings leaves model behavior effectively unchanged.
We also considered a similar low variance setup under the setting of Section 4.3. To do so, we modify the zero variance concatenated features with added mean-zero Gaussian noise in the training data. However, as we noted in the main paper, we found that unless the scaling of the added Gaussian noise was significantly less than $\gamma$ (i.e. $\gamma*10^{-3}$), the label smoothing and Mixup overfitting phenomenon was also exhibited by weight decay. This is possibly just a consequence of our experimental setup and/or the selected hyperparameters (i.e. perhaps this sensitivity is avoidable with some optimal choice of the weight decay hyperparameter), but we do not exhaustively explore this and instead leave a more comprehensive understanding of this phenomenon to follow-up work.

### v2

@@
 In this work, we analyze the role played by the label augmentation aspect of such methods.
-We prove that linear models on linearly separable data trained with label augmentation learn only the minimum variance features in the data, while standard training (which includes weight decay) can learn higher variance features.
-An important consequence of our results is negative: label smoothing and Mixup can be less robust to adversarial perturbations of the training data when compared to standard training.
-We verify that our theory reflects practice via a range of experiments on synthetic data and image classification benchmarks.
+We first prove that linear models on binary classification data trained with label augmentation learn only the minimum variance features in the data, while standard training (which includes weight decay) can learn higher variance features.
+We then use our techniques to show that even for nonlinear models and general data distributions, the label smoothing and Mixup losses are lower bounded by a function of the model output variance.
+An important consequence of our results is negative: label smoothing and Mixup can be less robust to spurious correlations in the data.
+We verify that our theory reflects practice via experiments on image classification benchmarks modified to have spurious correlations.
 1 Introduction
@@
 As a result, Mixup modifies not only the training labels but also the training inputs.
-The general principles of label smoothing and Mixup have been extended to create several variants, such as Structural Label Smoothing (Li et al., 2020), Adaptive Label Smoothing (Wang et al., 2021), Manifold Mixup (Verma et al., 2019), CutMix (Yun et al., 2019), PuzzleMix (Kim et al., 2020), SaliencyMix (Uddin et al., 2020), AutoMix (Liu et al., 2021), and Noisy Feature Mixup (Lim et al., 2022).
+The general principles of label smoothing and Mixup have been extended to several variants, such as Structural Label Smoothing (Li et al., 2020), Adaptive Label Smoothing (Wang et al., 2021), Manifold Mixup (Verma et al., 2019), CutMix (Yun et al., 2019), PuzzleMix (Kim et al., 2020), SaliencyMix (Uddin et al., 2020), AutoMix (Liu et al., 2021), and Noisy Feature Mixup (Lim et al., 2022).
 Due to the success of label smoothing and Mixup-based approaches, an important question on the theoretical side has been understanding when and why these data augmentations improve model performance.
@@
 Our analysis allows us to identify a clear failure mode shared by both Mixup and label smoothing - that they can ignore higher variance features - which we also verify empirically.
-1.1 Outline of Main Contributions
-Our main contribution can be summarized as,
-For linearly separable data distributions that can be learned using any one of multiple features, models trained with Mixup or label smoothing learn only the minimum variance feature.
-In order to work up to this claim, we provide the necessary background on Mixup and label smoothing and formalize the notion of data distributions with multiple separating features in Section 2.
-Our definition draws heavily from the idea of multi-view data introduced by Allen-Zhu & Li (2021), but is considerably simpler as it does not involve patches or feature noise.
-We then prove our main results in Section 3, where we show that linear models trained with Mixup or label smoothing on the aforementioned type of binary class data distribution are correlated only with the minimum variance feature in the data.
-On the other hand, we also show that linear models trained using just empirical risk minimization (ERM) and weight decay (so as to ensure a unique solution) have non-trivial correlation with the high variance feature in the data.
-A key aspect of our analysis is that we can show a clear separation in terms of feature learning between Mixup/label smoothing and ERM in the setting of just linear models and linearly separable data distributions, whereas previous feature learning analyses of Mixup/similar (Chidambaram et al., 2023; Zou et al., 2023) have relied on data distributions that are explicitly constructed to not be learnable by linear models and have an asymptotically large number of classes.
-As a result, our analysis is both simpler than existing analyses and shows that nonlinear training dynamics are not necessary to separate Mixup/label smoothing and ERM with respect to feature learning.
-We also verify that our theoretical results extend to practice across several experimental settings.
-In Section 4.1, we show that for synthetic data following our theoretical assumptions, logistic regression trained with Mixup or label smoothing learns a weight vector that is effectively uncorrelated with the high variance feature in the data.
-We also show that this behavior extends beyond synthetic data in Section 4.2, where we train logistic regression models on binary classification versions of CIFAR-10 and CIFAR-100 modified to have adversarial pixels in the training data that completely identify each class.
-In this case, the Mixup and label smoothing logistic regression models do not generalize at all due to having learned only the adversarial pixels in the training data.
-Finally, we show in Section 4.3 that these observations still apply even when we consider ResNets trained on the full versions of CIFAR-10 and CIFAR-100; for these experiments we follow the experimental setup introduced by Chidambaram et al.
-(2023) in which the training data is modified to have extra low variance features via additional channels.
-Our results provide a contrasting perspective to that of Chidambaram et al.
-(2023) - Mixup (and label smoothing) can end up learning only the low variance features in the training data, whereas ERM can avoid this problem.
+1.1 Main Contributions
+Our main contributions are three-fold:
+  1. 1.
+    We introduce a simple class of data distributions in which there exists a low variance feature and a high variance feature, and show that optimizing the label smoothing and Mixup losses for a linear model for such distributions leads to learning only the low variance feature.
+On the other hand, if the high variance feature satisfies a greater separability condition relative to the low variance feature, optimizing the empirical risk with weight decay successfully learns both features in the data.
+Our label smoothing and Mixup results are stated under general conditions on both the data and the respective loss hyperparameters; our Mixup results apply to any mixing distribution that does not coincide with ERM, which greatly generalizes the settings considered by prior work on Mixup feature learning (Chidambaram et al., 2023; Zou et al., 2023).
+  2. 2.
+    We show that our techniques extend to the analysis of general model classes, and prove lower bounds for the label smoothing and Mixup losses in terms of the model output variance.
+These results suggest that optimizing these losses necessarily decreases model output variance, which is not necessarily the case for training with weight decay.
+  3. 3.
+    We empirically verify our theoretical insights by considering image classification benchmarks modified to (1) have adversarial spurious correlations in the training data and (2) more natural spurious correlations in the form of class-correlated backgrounds.
+In both cases, we observe that label smoothing and Mixup fail to generalize for the full range of hyperparameters used in practice, while ERM with weight decay does generalize.
 1.2 Related Work
@@
 Mixup has also been studied from the perspective of calibration (Thulasidasan et al., 2019), with theoretical results (Zhang et al., 2021; Chidambaram & Ge, 2023) showing that Mixup training can prevent miscalibration issues arising from ERM (even when post-training calibration methods are used with ERM).
-Very recently, Oh & Yun (2023) studied Mixup in a similar context to our work (binary classification with linear models), and showed that Mixup can significantly improve the sample complexity required to learn the optimal classifier when compared to ERM.
-Also recently, and most closely related with our results in this paper, Mixup has been studied from a feature learning perspective, with Chidambaram et al.
+Recently, Oh & Yun (2023) studied Mixup in a similar context to our work (binary classification with linear models), and showed that Mixup can significantly improve the sample complexity required to learn the optimal classifier when compared to ERM.
+Most closely related with our results in this paper, Mixup has been studied from a feature learning perspective, with Chidambaram et al.
 (2023) and Zou et al.
@@
 2 Preliminaries
-Notation.
-Given $n\in\mathbb{N}$, we use $[n]$ to denote the set $\{1,2,...,n\}$.
-For a vector $x\in\mathbb{R}^{d}$ and a subset $S\subset[d]$, we use $x_{S}\in\mathbb{R}^{\absolutevalue{S}}$ to denote the restriction of $x$ to only those indices in $S$, and also use $x_{i}$ to denote the $i^{\text{th}}$ coordinate of $x$.
-The same notation also applies to matrices; i.e.
-for a matrix $T$ we use $T_{S}$ to denote the restriction of the rows of $T$ to only those dimensions in $S$.
-Additionally, we use $\succ$ to denote the partial order over positive definite matrices and $\mathrm{Id}$ to denote the identity matrix.
-We use $\norm{\cdot}$ to indicate the Euclidean norm on $\mathbb{R}^{d}$.
-For two sets of points $A,B\subset\mathbb{R}^{d}$ we use $d(A,B):=\min_{x\in A,y\in B}\|x-y\|$ to denote the distance between $A$ and $B$.
-For a function $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ we use $g^{i}(x)$ to denote the $i^{\text{th}}$ coordinate function of $g$.
-For a probability distribution $\pi$ we use $\mathrm{supp}(\pi)$ to denote its support.
-Additionally, if $\pi$ corresponds to the joint distribution of two random variables $X$ and $Y$ (i.e.
-data and label), we use $\pi_{X}$ and $\pi_{Y}$ to denote the respective marginals, and $\pi_{X\mid Y=y}$ and $\pi_{Y\mid X=x}$ to denote the conditional distributions.
-Our theory will focus on binary classification problems in $\mathbb{R}^{d}$ in which a subset of the input dimensions correspond to a low variance feature, and the complementary dimensions correspond to a higher variance but more separable feature.
-This setup is intended to mimic standard explorations of robustness of image classification models.
-For example, if each class in the data has a small set of constant, identifying pixels in every training point (either introduced adversarially or as an artifact of the domain being considered), does a model trained on this data learn to predict using only the identifying pixels or the whole image?
-The goal in such cases is typically to also learn the features present in the non-identifying pixels (i.e.
-the higher variance features).
-We will show, however, that doing any kind of label augmentation (label smoothing or Mixup) in this low variance/high variance feature setup will lead to learning a model that has only learned the low variance feature, while standard training will lead to having non-trivial correlation with the high variance feature.
-It is also natural to ask whether this phenomenon is purely a result of the fact that augmented labels (i.e.
-not 0 or 1) prevent the learned model weights from going to infinity, which has been pointed out to be crucial in previous existing analyses of Mixup (Chidambaram et al., 2023).
-We answer this question in the negative by also showing that training using ERM with weight decay still learns the high variance feature in the data, so the variance minimization effect is more than just a weight norm constraint.
-2.1 Data Model
-We begin with the generative model of the data that we will work with.
-For simplicity, we consider the low variance feature to have zero variance conditional on the label $y$ it is not difficult to extend our results to the case where the variance of this feature is simply much less than that of the high variance feature, but considering zero conditional variance yields cleaner proofs.
-We also consider balanced classes to further simplify the proofs, but our results will hold so long as each class appears with some non-zero probability.
+[…]
+We consider the $k$-class classification setting, in which there is a ground-truth data distribution $\pi$ on $\mathbb{R}^{d}\times[k]$ and our goal is to model the conditional distribution $\pi_{Y\mid X}$ using a learned function $g$.
+In our main theoretical results, we will pay particular attention to the case where $k=2$ and $g$ is a logistic regression model (although we generalize a weak version of these observations to $k$ classes as well), as in this setting we can get a clear handle on the features in the data learned by an optimal model with respect to a particular loss.
+For this case, we will assume that $\pi$ is supported on $\mathbb{R}^{d}\times\{\pm 1\}$ and that $g$ is parameterized by a weight vector $w$, i.e.
+$g_{w}(x)=\sigma(w^{\top}x)$, where $\sigma$ is the sigmoid function.
+Throughout this work, we will consider the following three families of losses.
+Standard cross-entropy with optional weight decay. The canonical cross-entropy (or negative log-likelihood) objective in the $k$-class setting is defined as:
+$\ell(g)=\mathbb{E}_{(X,Y)\sim\pi}[-\log g^{Y}(X)].$ (2.1)
+Here we have not specified a weight decay term, since we have placed no constraints on the structure of $g$.
+On the other hand, for linear binary classification, we can define the binary cross-entropy with optional weight decay as (recalling that $Y\in\{\pm 1\}$)
+$\ell_{\beta}(w)$ $=\mathbb{E}_{(X,Y)\sim\pi}\left[-\log g_{w}(YX)\right]+\frac{\beta}{2}\norm{w}^{2},$ (2.2)
+where we have defined the loss in terms of the parameter vector $w$.
+When $\beta>0$, (2.2) has a unique minimizer and we will directly analyze its properties.
+On the other hand, when $\beta=0$ (no weight decay), this is no longer the case - but our results will still apply to the common case of minimizing (2.2) by scaling the max-margin solution, which we define below.
 Definition 2.1.
-[Data Distribution] Let $\mathcal{L}\subset[d]$ and $\mathcal{H}=\mathcal{L}^{c}$ be the sets denoting the low variance feature and high variance feature in the data respectively.
-We define $\pi$ to be any distribution on $\mathbb{R}^{d}\times\{-1,1\}$ with $E_{x\sim\pi_{X}}[x]=0$ and $\pi_{Y}(y)=1/2$ for both $y\in\{-1,1\}$, along with the following properties:
-  1. 1.
-    (Zero Conditional Variance in $\mathcal{L}$) For $\pi_{X}$-a.e.
-$x\in\mathrm{supp}(\pi_{X\mid Y=y})$, we have that there exists a $v_{y}^{*}\in\mathbb{R}^{\absolutevalue{\mathcal{L}}}$ such that $x_{\mathcal{L}}=v_{y}^{*}$ with $\norm{v_{1}^{*}-v_{-1}^{*}}>0$.
-  2. 2.
-    (Non-zero Conditional Variance in $\mathcal{H}$) Let $\mu_{y}=\mathbb{E}[X\mid Y=y]$ and $\Sigma_{y}=\mathbb{E}[(x-\mu_{y})(x-\mu_{y})^{\top}\mid Y=y]$, then $\Sigma_{y,\mathcal{H}}$ (the restriction of $\Sigma_{y}$ to $\mathcal{H}$) is positive definite.
-  3. 3.
-    (More Separable High Variance Feature) Let $A_{y}=\{x_{\mathcal{H}}:x\in\mathrm{supp}(\pi_{X\mid Y=y})\}$.
-Then $A_{1},A_{-1}$ are compact, convex, and disjoint and satisfy $d(A_{1},A_{-1})>\norm{v_{1}^{*}-v_{-1}^{*}}$.
-Since Definition 2.1 corresponds to a linearly separable data distribution, we can restrict our attention to linear models of the form $g_{w}(x)=\sigma(w^{\top}x)$, where $\sigma(\cdot)$ is the sigmoid function.
-We can then define the binary cross-entropy with weight decay as:
-$\ell_{\lambda}(w)$ $=\mathbb{E}_{(x,y)\sim\pi}\left[-\log g_{w}(yx)\right]+\frac{\lambda}{2}\norm{w}^{2}$ (2.1)
-With $\ell_{0}(w)$ thereby indicating the regular binary cross-entropy without weight decay.
-In what follows, we will work with the population form of the loss as defined in Equation (2.1), since for $\lambda>0$ we have that $\ell_{\lambda}(w)$ is strictly convex and therefore we may simply analyze its unique minimizer as opposed to worrying about training dynamics.
-Although $\ell_{0}(w)$ will not have a unique minimizer, our results will still apply to the common case where $\ell_{0}(w)$ is minimized by scaling the max-margin solution, which we define below.
-Definition 2.2.
-[Max-Margin Solution] The max-margin solution $w^{*}$ with respect to $\pi$ is defined to be:
+[Max-Margin Solution] The max-margin solution $w^{*}$ with respect to $\pi$ is defined as:
 $w^{*}$ $=\operatorname*{argmin}_{w\in\mathbb{R}^{d}}\norm{w}^{2}$
 […]
-2.2 Label Smoothing
-Label smoothing as introduced by Szegedy et al.
-(2015) is intended to work with multi-class classification problems in which the data distribution $\pi$ is typically defined on $\mathbb{R}^{d}\times[k]$.
-In this case, we work with models $g:\mathbb{R}^{d}\to\Delta^{k}$ (the probability simplex in $\mathbb{R}^{k}$), and the cross-entropy loss is simply $\ell(g)=\mathbb{E}_{(x,y)\sim\pi}[-\log g^{y}(x)]$.
-The label-smoothed cross-entropy is then obtained by replacing the one-hot encoded label $y$ with a mixture of the original label and the uniform distribution over $[k]$.
+Label smoothing.
+The cross-entropy as defined in (2.1) treats the reference distribution that we compare $g(X)$ to as a point mass on the class $Y$.
+On the other hand, the label-smoothed cross-entropy is obtained by instead treating the reference distribution as a mixture of a point mass on $Y$ and the uniform distribution over $[k]$.
 Namely, the label-smoothed loss with mixing hyperparameter $\alpha\in[0,1]$ is defined to be:
-$\ell_{\mathrm{LS},\alpha}(g)=-\mathbb{E}_{(x,y)\sim\pi}\bigg{[}$ $(1-\alpha)\log g^{y}(x)+\frac{\alpha}{k}\sum_{i=1}^{k}\log g^{i}(x)\bigg{]}$ (2.3)
-For the purposes of our theory, however, we will be working with models of the form $g_{w}(x)=\sigma(w^{\top}x)$ and labels $y\in\{-1,1\}$, so we will consider the binary label-smoothed loss defined as:
-$\ell_{\mathrm{LS},\alpha}(w)=-\mathbb{E}_{(x,y)\sim\pi}\bigg{[}$ $\left(1-\frac{\alpha}{2}\right)\log g_{w}(yx)+\frac{\alpha}{2}\log g_{w}(-yx)\bigg{]}$ (2.4)
-Note that similar to Equation (2.1) the latter is defined as a function of the weight vector $w$ instead of the full model $g$.
-2.3 Mixup
-Similar to label smoothing, Mixup (Zhang et al., 2018) was also introduced in the context of multi-class classification.
-In addition to augmenting the one-hot encoded labels $y$, Mixup also augments the input data $x$ by combining both with another data point $(x^{\prime},y^{\prime})$ via a random convex combination.
-Both of these transformations are done according to a mixing distribution $\mathcal{D}_{\lambda}$ (which is a hyperparameter) whose support is contained in $[0,1]$.
-To simplify notation, we will use $x_{1:n}$ and $y_{1:n}$ to denote multiple data points $x_{1},...,x_{n}$ and their corresponding labels $y_{1},...,y_{n}$.
-Additionally, in the multi-class setting with $g:\mathbb{R}^{d}\to\Delta^{k}$, we also define:
-$h_{\lambda}(g,x_{1:2},y_{1:2})$ $=\lambda\log g^{y_{1}}(z_{\lambda})+(1-\lambda)\log g^{y_{2}}(z_{\lambda})$ (2.5)
-$\text{where}\quad z_{\lambda}$ $=\lambda x_{1}+(1-\lambda)x_{2}$ (2.6)
-We may then define the Mixup cross-entropy:
-$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(g)=\mathbb{E}_{(x_{1:2},y_{1:2})\sim\pi,\lambda\sim\mathcal{D}_{\lambda}}\left[-h_{\lambda}(g,x_{1:2},y_{1:2})\right]$ (2.7)
-Where $(x_{1:2},y_{1:2})\sim\pi$ is shorthand for $(x_{1},y_{1}),(x_{2},y_{2})\sim\pi$.
-As before, since our theory focuses on binary classification and linear models $g_{w}$, we may identically define the binary version $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ of Equation (2.7) except for redefining $h_{\lambda}$ to be:
-$h_{\lambda}(w,x_{1:2},y_{1:2})$ $=\lambda\log g_{w}(y_{1}z_{\lambda})+(1-\lambda)\log g_{w}(y_{2}z_{\lambda})$ (2.8)
-Again, observe that we have redefined $h_{\lambda}$ to directly be a function of $w$ instead of $g$.
+$\ell_{\mathrm{LS},\alpha}(g)=-\mathbb{E}_{(X,Y)\sim\pi}\bigg{[}(1-\alpha)\log g^{Y}(X)+\frac{\alpha}{k}\sum_{i=1}^{k}\log g^{i}(X)\bigg{]}.$ (2.4)
+And the corresponding binary version is (once again, $Y\in\{\pm 1\}$):
+$\ell_{\mathrm{LS},\alpha}(w)=-\mathbb{E}_{(X,Y)\sim\pi}\bigg{[}$ $\left(1-\frac{\alpha}{2}\right)\log g_{w}(YX)+\frac{\alpha}{2}\log g_{w}(-YX)\bigg{]}.$ (2.5)
+Mixup.
+Similar to label smoothing, Mixup also augments the reference distribution from being a point mass on $Y$ to being a mixture.
+However, Mixup also augments the input data as well.
+In particular, Mixup considers convex combinations of two pairs of points $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$, with a mixing weight sampled from a distribution $\mathcal{D}_{\lambda}$ (which is a hyperparameter) whose support is contained in $[0,1]$.
+To simplify notation, we will use $X_{1:n}$ and $Y_{1:n}$ to denote multiple inputs $X_{1},...,X_{n}$ and their corresponding labels $Y_{1},...,Y_{n}$, and additionally introduce a function $h$ defined as:
+$h(\lambda,g,X_{1:2},Y_{1:2})$ $=\lambda\log g^{Y_{1}}(Z_{\lambda})+(1-\lambda)\log g^{Y_{2}}(Z_{\lambda})$ (2.6)
+$\text{where}\quad Z_{\lambda}$ $=\lambda X_{1}+(1-\lambda)X_{2}.$ (2.7)
+After which we can define the Mixup cross-entropy as:
+$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(g)=\mathbb{E}_{(X_{1:2},Y_{1:2})\sim\pi,\lambda\sim\mathcal{D}_{\lambda}}\left[-h(\lambda,g,X_{1:2},Y_{1:2})\right].$ (2.8)
+The corresponding binary version $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ of (2.8) is identical except for redefining $h$ to be:
+$h(\lambda,w,x_{1:2},y_{1:2})$ $=\lambda\log g_{w}(Y_{1}Z_{\lambda})+(1-\lambda)\log g_{w}(Y_{2}Z_{\lambda}).$ (2.9)
 3 Main Theoretical Results
-Having established the necessary background, we now state our main results.
-Full proofs of all results can be found in Appendix A.
-We first show that the unique minimizer $w^{*}$ of the standard risk with weight decay (as defined in Equation (2.1)) has non-trivial correlation with the high variance dimensions of the input $x$.
+All omitted proofs in this section can be found in Section A of the Appendix.
+3.1 Linear Binary Classification
+We begin first with the binary setting, in which we will consider a data model where a subset of the input dimensions correspond to a low variance feature and the complementary dimensions correspond to a high variance feature that is more separable.
+This setup is intended to be a simplistic model of spurious correlations that appear in image data; one could think of the low variance feature as a fixed background while the high variance feature corresponds to the actual object category of interest.
+The goal in such cases is to ensure that we do not only learn the low variance feature, so that generalization is possible even under a shift of this feature (for example, recognizing the same object on different backgrounds).
+We will show, however, that doing any kind of label augmentation (label smoothing or Mixup) in this low variance/high variance feature setup will lead to a model that has only learned the low variance feature, whereas minimizing the binary cross-entropy with non-zero weight decay requires learning the high variance feature due to its greater separability.
+The overarching data model for this subsection is as follows.
+Definition 3.1.
+[Binary Data Distribution] We consider $\pi$ to be a distribution supported on $K\times\{\pm 1\}$ where $K$ is a compact subset of $\mathbb{R}^{d}$.
+We assume that $\pi$ is nondegenerate in that it satisfies $\pi_{Y}(y)>0$ for each $y\in\{\pm 1\}$ and that $\Sigma_{X}$ is positive definite, and we also assume $\mathbb{E}[X]=0$.
+Additionally, we designate a subset $\mathcal{L}\subseteq[d]$ that we refer to as the low variance feature in the data, and we refer to the complement $\mathcal{H}=\mathcal{L}^{c}$ as the high variance feature.
+For convenience, for a vector $v\in\mathbb{R}^{d}$, we will use $v\in\mathcal{L}$ to mean that only the $\mathcal{L}$ dimensions of $v$ are non-zero.
+Definition 3.1 does not actually make any strong assumptions on $\pi$ yet; this is because the weight decay and label augmentation results have different, non-conflicting assumptions that we codify separately.
+For the former, we need the data to be more separable in the $\mathcal{H}$ dimensions than in the $\mathcal{L}$ dimensions, in the following sense.
+Assumption 3.2.
+We assume that for every unit vector $u^{*}\in\mathcal{L}$, there exists a unit vector $v^{*}\in\mathcal{H}$ and $y\left\langle v^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$ for $\pi$-a.e.
+$(x,y)$.
+Assumption 3.2 is strong, but it actually does not imply linear separability, only that the $\mathcal{H}$ dimensions are in a sense better than the $\mathcal{L}$ dimensions for classification.
+For a simple example to provide intuition for the assumption, one can consider a 2-D distribution in which the $x$ coordinates for both classes are concentrated close to 0, but the $y$ coordinates are either large negative or large positive values depending on the class.
+Our first result shows that for $\ell_{\beta}(w)$ as defined in (2.2), the minimizer $w^{*}$ has a large correlation with the dimensions in $\mathcal{H}$.
+This is of course intuitive given Assumption 3.2, but it is not immediate because the weight decay penalty in (2.2) encourages distributing norm across all of the dimensions of $w$ (e.g.
+$\sum w_{i}$ is maximized with respect to the constraint $\norm{w}=1$ by considering $w_{i}=1/\sqrt{d}$).
 Theorem 3.3.
-Let $w^{*}$ be the unique minimizer of $\ell_{\lambda}(w)$ for $\lambda>0$.
-Then $\norm{w^{*}_{\mathcal{H}}}^{2}\geq\frac{1}{2}\norm{w^{*}}^{2}$.
-In other words, the optimal solution under weight decay has significant weight in the high variance dimensions $\mathcal{H}$.
+Let $w^{*}$ be the unique minimizer of $\ell_{\beta}(w)$ for $\beta>0$ under $\pi$ satisfying Assumption 3.2. Then $\norm{w^{*}_{\mathcal{H}}}^{2}\geq\frac{1}{2}\norm{w^{*}}^{2}$.
 Proof Sketch.
-We can orthogonally decompose the optimal solution $w^{*}$ in terms of the unit normal max-margin directions $u_{1}^{*}$ and $u_{2}^{*}$ separating $x_{\mathcal{L}}$ and $x_{\mathcal{H}}$ respectively.
-Then, using the fact that $y\left\langle u_{2}^{*},x_{\mathcal{H}}\right\rangle>y\left\langle u_{1}^{*},x_{\mathcal{L}}\right\rangle$ for all $(x,y)\in\mathrm{supp}(\pi)$ (from assumption three in Definition 2.1), we can claim that $w^{*}$ must have greater weight associated with $u_{2}^{*}$ than $u_{1}^{*}$, as otherwise we can decrease $\ell_{\lambda}(w^{*})$ by moving some weight from $u_{1}^{*}$ to $u_{2}^{*}$.
+We can orthogonally decompose the optimal solution $w^{*}$ in terms of unit normal directions $u^{*}\in\mathcal{L}$ and $v^{*}\in\mathcal{H}$.
+We show that in this decomposition it must be the case that $y\left\langle v^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$ for $\pi$-a.e.
+$(x,y)$, and then we can claim that $w^{*}$ must have greater weight associated with $v^{*}$ than $u^{*}$, as otherwise we can decrease $\ell_{\beta}(w^{*})$ by moving weight from $u^{*}$ to $v^{*}$.
 We cannot immediately extend the result of Theorem 3.3 to the binary cross-entropy $\ell_{0}(w)$ without weight decay, since there is no unique minimizer of $\ell_{0}$.
@@
 If $w^{*}$ is the max-margin solution to $\ell_{0}(w)$, then the result of Theorem 3.3 still holds.
-On the other hand, as alluded to earlier, this phenomenon does not occur when considering the label smoothing and Mixup losses $\ell_{\mathrm{LS},\alpha}(w)$ and $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$.
-Indeed, we can show:
+On the other hand, as alluded to earlier, this phenomenon does not occur when considering the label smoothing and Mixup losses $\ell_{\mathrm{LS},\alpha}$ and $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$.
+We will show that for both label smoothing and Mixup, the optimal solutions have arbitrarily small correlation with the dimensions in $\mathcal{H}$ as the distribution of $YX$ concentrates.
+For these results, we require the following assumptions (but no longer need Assumption 3.2).
+Assumption 3.5.
+We assume that $\mathbb{E}[YX_{\mathcal{L}}]\neq 0$ and $\Sigma_{YX,\mathcal{H}}\succ\rho\mathrm{Id}$ for some $\rho>0$.
+Here $\Sigma_{YX,\mathcal{H}}$ denotes the covariance matrix of $\Sigma_{YX}$ restricted to those rows and columns in $\mathcal{H}$.
+The first part of Assumption 3.5 just ensures it is possible to obtain a good solution using the dimensions in $\mathcal{L}$ while the second part codifies the idea of $\mathcal{H}$ being a high (at least non-zero) variance feature.
+Observe that we have made no direct separability assumptions on the data, although it is true that as $\norm{\Sigma_{YX,\mathcal{L}}}\to 0$ the class-conditional supports are guaranteed to be linearly separable in $\mathcal{L}$.
+We now state both results.
 Theorem 3.6.
-Every minimizer $w^{*}$ of $\ell_{\mathrm{LS},\alpha}(w)$ for $\alpha>0$ satisfies $\norm{w^{*}_{\mathcal{H}}}=0$, so the optimal solutions under label smoothing have no weight in the high variance dimensions.
+For $\alpha\in(0,1)$ and $\pi$ satisfying Assumption 3.5, every minimizer $w^{*}$ of $\ell_{\mathrm{LS},\alpha}(w)$ satisfies $\norm{w^{*}_{\mathcal{H}}}<O(\norm{\Sigma_{YX,\mathcal{L}}})$.
 Proof Sketch.
-The optimal prediction at a label-smoothed point is either $21-\alpha/$ 2 or $\alpha/2$ (depending on whether the label is $1$ or $-1$ respectively).
-We can thus construct a predictor that is optimal at every point by choosing $w$ to have the appropriate correlation with $v_{1}^{*}$ and having no correlation with $x_{\mathcal{H}}$.
-That there are no equally good predictors that have correlation with $x_{\mathcal{H}}$ can be shown by appealing to convexity.
-Analogously, we have the following result for Mixup:
+We show that $\ell_{\mathrm{LS},\alpha}(w)$ is strongly convex in $w^{\top}YX$.
+We then use Jensen’s inequality to get a lower bound on the loss in terms of this quantity, and show that this lower bound is achievable using only $w\in\mathcal{L}$ as the variance of $YX_{\mathcal{L}}$ decreases.
+Then using a lower bound on the Jensen gap (Lemma A.1) and Assumption 3.5, we show that any solution that is sufficiently non-zero in the $\mathcal{H}$ dimensions remains bounded away from this lower bound.
 Theorem 3.7.
-Every minimizer $w^{*}$ of $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ satisfies $\norm{w^{*}_{\mathcal{H}}}=0$ for any distribution $\mathcal{D}_{\lambda}$ with $\mathrm{supp}(\mathcal{D}_{\lambda})\subseteq[0,1]$ that is not a point mass on either 0 or 1.
+For any symmetric $\mathcal{D}_{\lambda}$ that is not a point mass on 00 or $1$ and $\pi$ satisfying Assumption 3.5, every minimizer $w^{*}$ of $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ satisfies $\norm{w^{*}_{\mathcal{H}}}<O(\norm{\Sigma_{YX,\mathcal{L}}})$.
 Proof Sketch.
-The proof setup is similar to that of Theorem 3.3, i.e.
-we try to find a stationary point of $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$ that is zero in the high variance dimensions.
-The difficulty in the Mixup analysis is that it is not possible to directly construct a point-wise optimal predictor like in the proof of Theorem 3.3 due to the different possibilities for mixed points (and mixed labels), and consequently there is no obvious construction for a stationary point.
-However, we can analyze the limiting behavior of the Mixup gradient with respect to the weight correlation with the low variance dimensions and still show that such a stationary point must exist.
+The proof follows the same outline as the label smoothing result, but there are now additional difficulties due to mixing inputs.
+In particular, we need to condition on $\lambda$ and then show strong convexity of the conditional loss in terms of $w^{\top}Z_{\lambda}$.
+We can then get a lower bound on the overall loss as an expectation with respect to $\lambda$ of this conditional loss.
+However, it is no longer immediate that there exists a stationary point $w\in\mathcal{L}$ that minimizes this lower bound.
+We prove the existence of such a stationary point by considering limiting behavior of the gradient of the loss as $w_{\mathcal{L}}$ tends to $\infty$ or $\infty$- ∞ in each component.
+After establishing a stationary point, we can show as with the label smoothing result that as $\norm{\Sigma_{YX,\mathcal{L}}}\to 0$ we can get a solution with $w\in\mathcal{L}$ that achieves the lower bound while solutions with $w_{\mathcal{H}}$ non-zero are bounded away from the lower bound.
+3.2 General Multi-Class Classification
+Our results so far have demonstrated a separation between standard training (ERM + weight decay) and label augmentation (label smoothing and Mixup) in the linear binary classification setting, without explicitly having to assume linear separability.
+The benefit of the linear binary case is that in this case the learning problems are convex in the weight vector $w$, so we can directly discuss properties of the optimal solutions instead of worrying about optimization dynamics.
+Of course, this is no longer true when we pass to nonlinear models, and standard model choices (e.g.
+neural networks) make it so that it is no longer easy to prove that a model has “learned” either $X_{\mathcal{L}}$ or $X_{\mathcal{H}}$ without explicitly analyzing some choice of optimization algorithm (i.e.
+gradient descent).
+However, our observations do translate to the outputs of any model.
+It is obvious that the model output variance should go to zero for any model that achieves the global optimum of the label augmentation losses we have discussed; indeed, this just corresponds to predicting the correct labels ($\alpha,1-\alpha$ with label smoothing and $\lambda,1-\lambda$ for every $\lambda$ for Mixup) with probability 1.
+On the other hand, it is less clear that we can make a quantitative statement regarding how bad the loss could be given a certain amount of variance in the model output.
+By lifting the techniques of the previous subsection, we can prove such quantitative results for both label smoothing and Mixup in the general multi-class setting.
+The general data distribution we consider for these results is as follows.
+Definition 3.8.
+[Multi-Class Data Distribution] We consider $\pi$ to be a distribution supported on $B\times[k]$ where $B$ is a compact subset of $\mathbb{R}^{d}$ and $k>2$.
+We assume only that $\pi$ satisfies $\pi_{Y}(y)>0$ for every $y$ and $\Sigma_{X}$ is positive definite.
+We make virtually no assumptions on $\pi$ for these results because we will only be proving general lower bounds, which are weaker than the claims regarding the optimal solutions of the previous subsection.
+For $\pi$ as in Definition 3.8, the label smoothing and Mixup results are as follows.
+Proposition 3.9.
+For $\alpha>0$ and any $g:\mathbb{R}^{d}\to\Delta^{k-1}$, letting $\mathrm{OPT}_{\mathrm{LS},\alpha}$ denote the minimum of $\ell_{\mathrm{LS},\alpha}$, we have for a universal constant $C>0$:
+$\ell_{\mathrm{LS},\alpha}(g)\geq\mathrm{OPT}_{\mathrm{LS},\alpha}+C\sum_{y=1}^{k}\pi_{Y}(y)\mathrm{Var}\left(g(X)\mid Y=y\right).$ (3.1)
+Proposition 3.10.
+For any $\mathcal{D}_{\lambda}$ that is not a point mass on 00 or $1$ and any $g:\mathbb{R}^{d}\to\Delta^{k-1}$, letting $\mathrm{OPT}_{\mathrm{MIX},\mathcal{D}_{\lambda}}$ denote the minimum of $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$, we have for a universal constant $C>0$ that:
+$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(g)\geq\mathrm{OPT}_{\mathrm{MIX},\mathcal{D}_{\lambda}}+C\sum_{y_{1}=1}^{k}\sum_{y_{2}=1}^{k}\pi_{Y}(y_{1})\pi_{Y}(y_{2})\mathbb{E}_{\lambda}\left[\mathrm{Var}\left(g(Z_{\lambda})\mid y_{1},y_{2},\lambda\right)\right].$ (3.2)
+The proofs for both Propositions 3.9 and 3.10 follow the same structure; we show that after appropriate conditioning, both losses can be broken up into a sum of strongly convex conditional losses.
+The lower bounds in both results show that in order to make progress with respect to the label smoothing or Mixup losses, a training algorithm needs to push model outputs to be low variance.
+Remark 3.11.
+We do not prove an analogous result to Theorem 3.3 in the general setting because Propositions 3.9 and 3.10 operate directly on the model outputs $g(X)$, and it is not clear how to translate a weight norm constraint to this setting without explicitly parameterizing $g$.
+Intuitively, however, with a weight norm constraint on $g$ it may no longer be optimal to have zero variance predictions, since such predictions may require very large weights depending on the data distribution.
 Comparisons to existing results.
 We are not aware of any existing results on feature learning for label smoothing; prior theoretical work largely focuses on the relationship between label smoothing and learning under label noise, which is orthogonal to the perspective we take in this paper.
-Although feature learning results exist for Mixup (Chidambaram et al., 2023; Zou et al., 2023), the pre-existing results are constrained to the case of mixing using point mass distributions and only prove a separation between Mixup and unregularized ERM, whereas our results work for arbitrary mixing distributions (that don’t coincide with ERM) and also separate Mixup from ERM with weight decay.
-This greater generality comes at the cost of considering only linear models and linearly separable data, however, as both Chidambaram et al.
+Although feature learning results exist for Mixup (Chidambaram et al., 2023; Zou et al., 2023), the pre-existing results are constrained to the case of mixing using point mass distributions and only prove a separation between Mixup and unregularized ERM, whereas our results work for arbitrary symmetric mixing distributions (that don’t coincide with ERM) and also separate Mixup from ERM with weight decay.
+This greater generality comes at the cost of considering only linear models for our feature learning results; both Chidambaram et al.
 (2023) and Zou et al.
-(2023) consider non-linearly separable data and the training dynamics of 2-layer neural networks.
+(2023) consider the training dynamics of 2-layer neural networks on non-separable data.
 Additionally, our results can be viewed as generalizing the observations made by Chidambaram et al.
 (2023), which were that Mixup can learn multiple features in the data when doing so decreases the variance of the learned predictor.
-In our case, we directly show that Mixup will learn the low variance feature in the data as opposed to the high variance feature.
+In our case, we directly show that Mixup will have much larger correlation with the low variance feature in the data as opposed to the high variance feature.
 Our results also do not contradict the observations of Zou et al.
@@
 4 Experiments
-(a) Weight Decay
-(b) Label Smoothing
-(c) Mixup
-Figure 1: Final weight norm of logistic regression model in the high variance dimensions ($\norm{w_{\mathcal{H}}}$) for various hyperparameter settings on the synthetic data distribution of Section 4.1.
-We now verify the predictions of our theoretical results in Section 3 across a variety of experimental setups.
+We now verify the predictions of our theoretical results in Section 3 across both binary and multi-class experimental setups.
 All experiments in this section were conducted on a single P100 GPU using PyTorch (Paszke et al., 2019) for model implementation.
 All reported results correspond to means over 5 training runs, and shaded regions in the figures correspond to 1 standard deviation bounds.
-4.1 Synthetic Data
-We first consider the case of training logistic regression models on a synthetic data distribution that exactly follows the assumptions of Definition 2.1.
-Definition 4.1 (Synthetic Data).
-We define a distribution $\pi_{\gamma}$ parameterized by $\gamma$ (where $0<\gamma\ll 1$) on $\mathbb{R}^{d}\times\{-1,1\}$ with $\pi_{\gamma,Y}(1)=1/2$ and $x\sim\pi_{\gamma,X\mid Y=y}$ satisfying:
-  1. 1.
-    First $\lfloor d/2\rfloor$ dimensions are constant but small. We have $x_{i}=\gamma y$ for $i\leq\lfloor d/2\rfloor$.
-  2. 2.
-    Last $d-\lfloor d/2\rfloor$ dimensions are high variance. We have $x_{i}\sim\mathrm{Uniform}([y,100y])$ for $i>\lfloor d/2\rfloor$.
-In other words, we consider data where (conditional on the label) the first half of the dimensions are fixed (corresponding to $\mathcal{L}$) and the second half are i.i.d.
-high variance uniform (corresponding to $\mathcal{H}$).
-We sample $n=5000$ data points according to Definition 4.1 with $d=10$ and $\gamma=0.1$ and train logistic regression models across a range of settings for weight decay, label smoothing, and Mixup.
-[^1] The choice of $\gamma$ here and throughout the rest of this section is relatively arbitrary; we verified that our experimental phenomena are unchanged for other scales such as $10^{-2},10^{-3},\text{ and }10^{-4}$.
-Namely, we consider 20 uniformly spaced values in $[0,0.1]$ for the weight decay $\lambda$ parameter and in $[0,0.75]$ for the label smoothing $\alpha$ parameter, where the upper bound for the label smoothing parameter space is obtained from the experiments of Müller et al.
-(2020).
-For Mixup, we fix the mixing distribution to be the canonical choice of $\mathrm{Beta}(\alpha,\alpha)$ introduced by Zhang et al.
-(2018) and consider 20 uniformly spaced $\alpha$ values in $[0,8]$ (with $\alpha=0$ corresponding to ERM), which effectively covers the range of Mixup hyperparameter values used in practice.
-We train all models for 100 epochs using Adam with the standard hyperparameters of $\beta_{1}=0.9,\beta_{2}=0.999$, a learning rate of $5*10^{-3}$, and a batch size of 500.
-At the end of training, we compute $\norm{w_{\mathcal{H}}}$ (i.e.
-the norm of the weight vector in the last 5 dimensions) for each trained model.
-For each model setting, we report the mean and standard deviation of $\norm{w_{\mathcal{H}}}$ over 5 training runs in Figure 1.
-As can be seen from the results, the weight decay models always have non-trivial values for $\norm{w_{\mathcal{H}}}$ (even for large values of $\lambda$), whereas the label smoothing and Mixup models very quickly converge to a $\norm{w_{\mathcal{H}}}$ of effectively zero as their respective hyperparameters move away from the ERM regime ($\alpha=0$ in both cases).
-This matches the behavior predicted by Theorems 3.1 to 3.4.
-It is natural to ask to what extent these results depend on the fact that the low variance features in Definition 4.1 (the first half of the dimensions) actually have zero variance conditional on the label.
-As we mentioned in our theoretical results, it is sufficient to assume that the variance of the dimensions in $\mathcal{L}$ is simply much less than that of $\mathcal{H}$, and the fact that we consider zero variance is merely for simplifying purposes.
-We verify that this is true in practice by re-running the above experiments but modifying the dimensions in $\mathcal{L}$ to have non-zero variance by adding uniformly distributed noise on the scale of $\gamma/10$.
-The results remain unchanged; the corresponding plots can be found in Section B of the Appendix.
-4.2 Binary Image Classification
+4.1 Spurious Pixels in Binary Classification
 (a) Weight Decay
@@
 Figure 1: Logistic regression final test errors for various hyperparameter settings on the binary classification versions of CIFAR-10 and CIFAR-100 from Section 4.1.
-Although the experiments of Section 4.1 constitute a direct verification of our theory, in practice the low and high variance features are typically not as obviously partitioned as in Definition 4.1 and the metric we care about is usually generalization error as opposed to the strength of correlation with some part of the data.
-As such, we now relax our strict adherence to the assumptions of Definition 2.1 and consider binary classification on image data.
-To do so, we reduce the standard benchmarks of CIFAR-10 and CIFAR-100 (Krizhevsky, 2009) to binary classification tasks by maintaining only the first two classes (“airplane” and “automobile” for CIFAR-10, “apple” and “aquarium fish” for CIFAR-100) and assigning them the labels of -1 and 1 respectively.
+We first consider the case of logistic regression trained on data that has trivial spurious correlations that do not appear in the test data.
+For our binary classification task, we consider reductions of the standard CIFAR-10 and CIFAR-100 (Krizhevsky, 2009) benchmarks to binary classification by fixing two classes from each dataset as the positive and negative classes, and replacing their original labels with the labels $+1$ and $-1$ respectively.
 This leads to 10000 training examples and 2000 test examples in the case of CIFAR-10, and 1000 training examples and 200 test examples in the case of CIFAR-100.
-We preprocess the training data to have mean zero and variance one along every image channel (and apply the same transform to the test data), and then modify the training data such that the top left value (corresponding to the “R” channel of the top left pixel in the original image) in the tensor representation of the training input is replaced by $y\gamma$, where $\gamma=0.1$ as in Section 4.1.
-This ensures that the training data is separable in the first dimension of the data, although unlike in the case of Definition 4.1 we make no modifications to the other dimensions of the data to ensure separability outside the first dimension.
+Our experiments are not sensitive to the choice of binary reduction, and for the experiments in this section we fix the $-1$ class to be class 00 from the original data and the $+1$ class to be class $1$ from the original data (“airplane” and “automobile” for CIFAR-10, “apple” and “aquarium fish” for CIFAR-100).
+We preprocess the training data to have mean zero and variance one along every image channel (and apply the same transform to the test data).
+To introduce spurious correlations into the training data, we modify the data such that the first value in the tensor representation of each training input (corresponding to the “R” channel of the top left pixel in the original image) is replaced by $y\gamma$ with $\gamma=0.1$ (the choice of $\gamma$ here is not important as long as it is small, we verified the results for $\gamma=10^{-5}$ up to $\gamma=0.1$).
+This ensures that the training data is trivially linearly separable in the first dimension of the data, but learning this first dimension requires having larger weight norm due to the scaling by $\gamma$.
 We leave the test data unchanged - our goal is to determine whether models trained on the modified training data can learn more than just the single identifying dimension.
-We follow the exact same logistic regression training setup as in Section 4.1, and sweep over the same hyperparameter values for weight decay, label smoothing, and Mixup.
-We report the mean test error along with one standard deviation error bounds over five training runs for each model setting in Figure 2.
-The results mirror our observations in Section 4.1: both the label smoothing and Mixup models have high test error for all settings (in this case, ERM performs poorly as well) while weight decay achieves a significantly lower test error for all $\lambda>0$.
+We then train logistic regression models to convergence (100 epochs) on both the reduced CIFAR-10 and CIFAR-100 tasks across a range of settings for weight decay, label smoothing, and Mixup.
+We consider 20 uniformly spaced values in $[0,0.1]$ for the weight decay $\lambda$ parameter and in $[0,0.75]$ for the label smoothing $\alpha$ parameter, where the upper bound for the label smoothing parameter is obtained from the experiments of Müller et al.
+(2020).
+For Mixup, we fix the mixing distribution to be the canonical choice of $\mathrm{Beta}(\alpha,\alpha)$ introduced by Zhang et al.
+(2018) and consider 20 uniformly spaced $\alpha$ values in $[0,8]$ (with $\alpha=0$ corresponding to ERM), which effectively covers the range of Mixup hyperparameter values used in practice.
+We train all models using Adam with the hyperparameters of $\beta_{1}=0.9,\beta_{2}=0.999$, a learning rate of $5*10^{-3}$, and a batch size of 500.
+After training, we report the mean test error along with one standard deviation error bounds over five training runs for each model setting in Figure 1.
+As can be seen from the results, both the label smoothing and Mixup models have high test error for all settings while weight decay achieves a significantly lower test error for all $\lambda>0$.
 The higher variance in the reported CIFAR-100 test errors can almost certainly be attributed to the fact that our binary version of CIFAR-100 consists of only 1000 training points and 200 points.
@@
 Figure 2: Comparison of norm ratio between first dimension (synthetically modified in the training data) and remaining dimensions (left unchanged) for trained logistic regression weight vector for the 20 different hyperparameter settings of weight decay, label smoothing, and Mixup.
-Our results correspond once again to label smoothing and Mixup learning to use only the zero variance, identifying dimension of the input, as we expect from our theory.
-Indeed, we verify this in Figure 3 where we plot the ratio $\norm{w_{1}}/\norm{w_{[d]\setminus\{1\}}}$ (i.e.
-the ratio between the norm of the trained model weight vector in the first dimension and the remaining dimensions).
-We observe that for both CIFAR-10 and CIFAR-100, across all the non-ERM hyperparameter settings (refer to Section 4.1 for the range of hyperparameters) that both label smoothing and Mixup place much greater weight in the first dimension than weight decay.
-As with the experiments of Section 4.1, we again check that the results here do not rely on the fact that the identifying dimension introduced into the training data has zero variance conditional on the label.
-As before, introducing uniformly distributed noise on the scale of $\gamma/10$ to the identifying dimension does not change our results, and the corresponding plots can again be found in Section B of the Appendix.
-4.3 Multi-Class Image Classification
-(a) Weight Decay ($\lambda=0.01$)
-(b) Label Smoothing ($\alpha=0.1$)
-(c) Mixup ($\alpha=1$)
-Figure 4: ResNet-18 test error at the end of each epoch of training on the modified versions of CIFAR-10 and CIFAR-100 from Section 4.3, using fixed settings for weight decay, label smoothing, and Mixup.
-Thus far we have considered training logistic regression models to stay within the confines of our theory, but we now investigate whether our observations extend to the more practical setting of training nonlinear models (i.e.
-deep neural networks) on multi-class image classification data.
-As in Section 4.2, we consider CIFAR-10 and CIFAR-100, but this time maintain all of the original classes in both datasets.
-We once again preprocess the training data to have mean zero and variance one, and apply this transform to the test data as well.
-Unfortunately, unlike in Section 4.2, introducing a zero variance, identifying feature into each class in the training data would now require modifying a more significant portion of each image (for example, a $10\times 10$ patch for CIFAR-100 to identify each class).
-Figure 5: Visualization of data modification applied to CIFAR-10/CIFAR-100 in Section 4.3.
-So as to preserve most of the original image content, as well as to have a more direct comparison to prior work, we choose to adopt a similar experimental setup to that of Chidambaram et al.
+Our results correspond to label smoothing and Mixup learning to use only the spurious, identifying dimension of the training inputs, as we expect from our theory since this dimension has zero conditional variance.
+Indeed, we verify this fact empirically in Figure 2, where we plot the ratio $\norm{w_{1}}/\norm{w_{[d]\setminus\{1\}}}$ (i.e.
+the ratio between the norm of the trained model weight vector in the first dimension and the remaining dimensions) for each of the trained models.
+4.2 Background Shift in Multi-Class Classification
+(a) Shaded MNIST
+(b) Model Performance
+Figure 3: Final model test errors over our hyperparameter sweep for the shaded MNIST dataset of Section 4.2, alongside a visualization of samples from the dataset.
+Although the experiments of Section 4.1 cleanly verify our theory, the spurious correlation introduced in the training data is effectively adversarial in nature.
+We now consider the case of a spurious correlation of greater practical interest: backgrounds that are correlated with the correct class label.
+In particular, we consider a very similar setup to the one used by Arjovsky et al.
+(2020) to motivate the influential invariant risk minimization framework: namely, we construct a shaded version of the MNIST dataset in which the background pixels for each class are replaced with different low intensity colors corresponding to the class labels.
+Unlike Arjovsky et al.
+(2020), however, we maintain all 10 classes since our theory from Section 3.2 suggests that our observations should generalize to the multi-class setting.
+The colors used for each class background are permuted between the train and the test data, so that a model that learns to predict only using background pixels cannot achieve good test accuracy.
+The only constraints we place on the background colors are that they are distinct across classes and that their intensities (i.e.
+values in RGB space) are small (it suffices to consider values bounded by 16); this makes it difficult to visually distinguish the background colors (hence "shaded MNIST") even though they are obviously different in RGB space.
+The color intensity constraint is not necessary for the failure of label smoothing and Mixup, but is necessary for weight decay to succeed since it ensures that the higher variance feature (the actual digit in the foreground) is generally more separable (due to larger pixel intensities).
+Note that although there is no variance in the background color conditional on a class, there is significant variance in the actual pixels per class as the locations of the background pixels change across data points.
+For our model setup, we also follow Arjovsky et al.
+(2020) and consider a simple 2-layer feedforward neural network with ReLU activations and a hidden layer size of $2048$, as this is sufficient to achieve good performance on MNIST and is a tiny enough model that we can efficiently do the same hyperparameter sweep from Section 4.1.
+We follow the same training procedure as Section 4.1, training until the training loss roughly converges (only 20 epochs are necessary).
+Test error results across the different hyperparameter settings for weight decay, label smoothing, and Mixup are shown in Figure 3.
+We observe the same phenomena as before, even in this non-trivial spurious correlation setting: both the label smoothing and Mixup models have high test error for all settings while weight decay achieves a significantly lower test error for all $\lambda>0$.
+5 Conclusion
+In this work, we have shown that label augmentation strategies such as label smoothing and Mixup exhibit a variance minimization effect that can lead to learning low variance, spurious correlations in the data even when there exists a better but higher variance feature.
+The main practical takeaway from this is that label smoothing and Mixup do not necessarily always improve the robustness of models, and their failure to do so may be indicative of certain identifying features in the training data (i.e.
+adversarially introduced patches or background pixels).
+Limitations.
+The main theoretical limitation of work is that our results for general models are only loss lower bounds and do not characterize optimization dynamics.
+Our experiments that such dynamics for feedforward networks would still reflect our observations, however, and it would be an interesting follow-up to study these dynamics using the techniques of Zou et al.
+(2023) and Chidambaram et al.
 (2023).
-Namely, we assign a random vector sampled from $\mathcal{N}(0,\gamma^{2}\mathrm{Id}_{d})$ (where $d$ is the dimension of the data) to each class and concatenate this class vector along the channels dimension to every training data point for that class.
-A visualization of this transformation is shown in Figure 5.
-For the test data points, we concatenate an all-zero vector of the same size so as to preserve the dimensionality across train and test, as was done by Chidambaram et al.
-(2023).
-We take $\gamma=0.1/\sqrt{d}$, so as to account for the dimensionality of the concatenated feature.
-For both datasets, we train ResNet-18 models (He et al., 2015) for 100 epochs using Adam with the standard hyperparameters of $\beta_{1}=0.9,\beta_{2}=0.999$, a learning rate of $10^{-3}$, and a batch size of 500.
-Due to the higher compute cost of training these models compared to the logistic regression models of Sections 4.1 and 4.2, we do not perform a full hyperparameter sweep but instead use known well-performing settings for weight decay, label smoothing, and Mixup.
-Namely, we use $\lambda=0.01$ for weight decay, $\alpha=0.1$ for label smoothing (Müller et al., 2020), and $\alpha=1$ in $\mathrm{Beta}(\alpha,\alpha)$ for Mixup (Zhang et al., 2018).
-Since we are now dealing with nonconvex training dynamics, we report the mean test error over five runs at the end of each epoch with one standard deviation error bounds for each model setting in Figure 4.
-As can be seen from the plots, both label smoothing and Mixup initially decrease the test error in the initial stages of training, but then overfit to the concatenated zero variance features.
-On the other hand, the test error for weight decay decreases over the course of training until effectively converging.
-This once again does not contradict what was observed by Chidambaram et al.
-(2023), as in their case their concatenated features were weighted alongside with the original image data so that variance could be decreased by learning both.
-We also considered the case of low variance concatenated features as opposed to zero variance concatenated features in the training data.
-Similar to the results of Sections 4.1 and 4.2, the label smoothing and Mixup overfitting phenomenons persist even when we introduce variance to the concatenated features.
-However, we found that unless the introduced variance is much smaller (i.e.
-multiple orders of magnitude) than the concatenated feature scale $\gamma$, weight decay also overfits to the concatenated features.
-This brittleness is possibly an artifact of the chosen experimental setup and/or hyperparameters, but due to compute constraints we did not exhaustively explore this behavior and instead leave it as an open direction for follow-up work extending our results to non-linear/non-convex settings.
-5 Conclusion
-In summary, we have shown that for linearly separable data distributions in which there exist high variance features and much lower variance features, linear models trained with the data augmentation schemes of label smoothing and Mixup fail to learn the low variance features in both practice and theory, while standard training with weight decay succeeds.
-The main practical takeaway from this is that label smoothing and Mixup do not necessarily always improve the robustness of models, and their failure to do so may be indicative of certain identifying features in the training data (i.e.
-adversarially introduced patches).
-A natural follow-up to the results we have presented is an extension to the case of non-linear models and non-linearly-separable data, which the results of Section 4.3 indicate should be possible.
+On the empirical side, our main limitation is that we have demonstrated our theory using only small-scale experiments; a comprehensive empirical evaluation of label smoothing and Mixup in larger scale spurious correlations settings guided by our theory would be a natural avenue for future work.
+Broader Impacts.
+Although label augmentation are used to train and finetune large-scale models, our results concerning them in this work have mostly been theoretical.
+As a result, we do not anticipate any direct misuse of our results or any broader harmful impacts.
 Acknowledgements
 Rong Ge and Muthu Chidambaram are supported by NSF Award DMS-2031849 and CCF-1845171 (CAREER). Muthu would like to thank Annie Tang for thoughtful feedback during the early stages of this project.
-Appendix A Full Proofs of Main Results
-Here we provide full proofs of all of the results contained in Section 3 of the main paper.
-A.1 Proof of Theorem 3.1 and Corollary 3.2
-See 3.1
+Appendix A Omitted Proofs
+A.1 Helper Lemma
+The following lemma lower bounds the Jensen gap in terms of the variance of the random variable being considered.
+We will make repeated use of it in the proofs of the label smoothing and Mixup results.
+Lemma A.1.
+Let $\phi:\mathbb{R}^{d}\to\mathbb{R}$ be a twice-differentiable convex function satisfying $\gamma_{1}\mathrm{Id}\prec\nabla^{2}\phi\prec\gamma_{2}\mathrm{Id}$.
+Then for any square integrable random variable $X$ on $\mathbb{R}^{d}$ it follows that:
+$\mathbb{E}[\phi(X)]-\phi(\mathbb{E}[X])\in\left[\frac{\gamma_{1}}{2}\mathrm{Var}(X),\frac{\gamma_{2}}{2}\mathrm{Var}(X)\right].$ (A.1)
 Proof.
-First, we claim that $w_{\mathcal{L}}^{*}$ is in the direction of $v_{1}^{*}$, which is the normal vector to the max-margin hyperplane separating $v_{1}^{*}$ and $v_{-1}^{*}$ (since $v_{1}^{*}=-v_{-1}^{*}$ due to the fact that the data distribution has mean zero and $\pi_{Y}(1)=\pi_{Y}(-1)$).
-Indeed, for any other choice of $w\in\mathbb{R}^{d}$ and any $(x,y)\in\mathrm{supp}(\pi)$, we have:
-$y\left\langle w_{\mathcal{L}},x_{\mathcal{L}}\right\rangle=y\left\langle w_{\mathcal{L}},v^{*}_{y}\right\rangle=\left\langle w_{\mathcal{L}},v^{*}_{1}\right\rangle$ (A.1)
-As a result, if $w_{\mathcal{L}}$ is not in the direction of $v_{1}^{*}$, we can remove the component orthogonal to $v_{1}^{*}$ (thereby decreasing the norm of the solution) while keeping the inner product with $x_{\mathcal{L}}$ the same.
-Now let $u_{1}^{*}=v_{1}^{*}/\norm{v_{1}^{*}}$ and $u_{2}^{*}\in\mathbb{R}^{\absolutevalue{\mathcal{H}}}$ similarly denote the unit normal vector to the max-margin hyperplane separating $A_{1}$ and $A_{-1}$.
-Recall that $\left\langle u_{2}^{*},x_{\mathcal{H}}\right\rangle>\left\langle u_{1}^{*},x_{\mathcal{L}}\right\rangle$ by assumption three in Definition 2.1.
-Let us identify $u_{1}^{*}$ and $u_{2}^{*}$ with the corresponding vectors in $\mathbb{R}^{d}$ whose $-\mathcal{L}$ and $-\mathcal{H}$ components are zero respectively, so that $u_{1}^{*}$ and $u_{2}^{*}$ are orthonormal in $\mathbb{R}^{d}$.
-Then we can decompose the unique minimizer of the loss $w^{*}$ as $\alpha u_{1}^{*}+\beta u_{2}^{*}+\kappa u_{3}^{*}$, where $u_{3}^{*}$ is a vector orthonormal to $u_{1}^{*}$ and $u_{2}^{*}$.
-Now suppose that $\alpha>\beta$. Then we claim that replacing $\alpha$ and $\beta$ with $\gamma=\sqrt{(\alpha^{2}+\beta^{2})/2}$ yields a solution with lower loss than $w^{*}$.
-To see this, first observe that $2\gamma^{2}=\alpha^{2}+\beta^{2}$, so that the norm of the modified solution is the same as $w^{*}$.
-This implies that the weight decay penalty term in $\ell_{\lambda}$ is unchanged.
-Furthermore, we have that $\gamma\in(\beta,\alpha)$, and that $\gamma-\beta>\alpha-\gamma$.
-This follows from the fact that $\gamma\geq(\alpha+\beta)/2$ by construction.
-This, combined with the fact that $y\left\langle u_{2}^{*},x_{\mathcal{H}}\right\rangle>y\left\langle u_{1}^{*},x_{\mathcal{L}}\right\rangle$, then implies for all $(x,y)\in\mathrm{supp}(\pi)$ that:
-$y\left\langle\gamma u_{1}^{*}+\gamma u_{2}^{*},x\right\rangle>y\left\langle\alpha u_{1}^{*}+\beta u_{2}^{*},x\right\rangle$ (A.2)
-Which contradicts the minimality of $w^{*}$. Therefore, we must have $\alpha\leq\beta$, which gives the desired result. ∎
-See 3.2
+From the assumption of the lemma it follows that $\phi(x)-\gamma_{1}\norm{x}^{2}/2$ is convex and $\phi(x)-\gamma_{2}\norm{x}^{2}/2$ is concave.
+Applying Jensen’s inequality to each of these functions yields (A.1).
+∎
+A.2 Proofs for Section 3.1
+See 3.3
+Proof.
+Let us decompose the unique minimizer $w^{*}$ as $w^{*}=\alpha_{1}u^{*}+\alpha_{2}v^{*}$, where $u^{*}$ and $v^{*}$ are orthonormal and satisfy $u^{*}_{\mathcal{H}}=0,v^{*}_{\mathcal{L}}=0$ (i.e.
+$u^{*}$ is a normalized version of the $\mathcal{L}$ components of $w^{*}$ and $v^{*}$ is the same but for the $\mathcal{H}$ components).
+We claim that $y\left\langle v^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$ for $\pi$-a.e.
+$(x,y)$.
+Indeed, if this were not the case then by the assumption on $\pi$ we could choose an orthonormal vector $z^{*}$ with $z^{*}_{\mathcal{L}}=0$ that satisfies $y\left\langle z^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$ for $\pi$-a.e.
+$(x,y)$ and decrease the loss of $w^{*}$ by replacing $v^{*}$ with $z^{*}$.
+Now suppose that $\alpha_{1}>\alpha_{2}$.
+Then we claim that replacing $\alpha_{1}$ and $\alpha_{2}$ with $\gamma=\sqrt{(\alpha_{1}^{2}+\alpha_{2}^{2})/2}$ yields a solution with lower loss than $w^{*}$.
+To see this, first observe that $2\gamma^{2}=\alpha_{1}^{2}+\alpha_{2}^{2}$, so that the norm of the modified solution is the same as $w^{*}$.
+This implies that the weight decay penalty term in $\ell_{\beta}$ is unchanged.
+Furthermore, we have that $\gamma\in(\alpha_{2},\alpha_{1})$, and that $\gamma-\alpha_{2}>\alpha_{1}-\gamma$.
+This follows from the fact that $\gamma\geq(\alpha_{1}+\alpha_{2})/2$ by construction.
+This, combined with the fact that $y\left\langle v^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$, then implies for $\pi$-a.e.
+$(x,y)$ that:
+$y\left\langle\gamma u^{*}+\gamma v^{*},x\right\rangle>y\left\langle\alpha_{1}u^{*}+\alpha_{2}v^{*},x\right\rangle.$ (A.2)
+Which contradicts the optimality of $w^{*}$. Therefore, we must have $\alpha_{1}\leq\alpha_{2}$, which gives the desired result. ∎
+See 3.4
 Proof.
@@
 ∎
-A.2 Proof of Theorem 3.3
-See 3.3
+See 3.6
 Proof.
-The strategy of the proof will be to construct a stationary point for the label smoothed loss that is correlated only with the dimensions in $\mathcal{L}$, and then argue using Jensen’s inequality that any solution that is correlated with the dimensions in $\mathcal{H}$ will have non-zero variance and thereby worse loss than our constructed stationary point.
-We first compute the gradient $\gradient_{w}\ell_{\mathrm{LS},\alpha}(w)$ and Hessian $\gradient_{w}^{2}\ell_{\mathrm{LS},\alpha}(w)$:
-$\gradient_{w}\ell_{\mathrm{LS},\alpha}(w)$ $=\mathbb{E}_{(x,y)\sim\pi}\left[(g_{w}(yx)-1+\alpha/2)yx\right]$ (A.3)
-$\gradient_{w}^{2}\ell_{\mathrm{LS},\alpha}(w)$ $=\mathbb{E}_{x\sim\pi_{X}}\left[g_{w}(yx)(1-g_{w}(yx))xx^{\top}\right]$ (A.4)
-Equations (A.3) and (A.4) follow from interchanging the gradient operator with the expectations, which is justified in this case via the dominated convergence theorem because $g_{w}(x)$ is continuously differentiable (with bounded derivative) in each component of $w$, and $x$ has bounded support.
-Clearly the Hessian is positive semidefinite, so $\ell_{\mathrm{LS},\alpha}(w)$ is convex.
-It is also straightforward to construct a stationary point $w^{*}$ of $\ell_{\mathrm{LS},\alpha}(w)$ by choosing $w^{*}_{\mathcal{L}}=\beta v_{1}^{*}$ for appropriate $\beta\in\mathbb{R}$ and $w^{*}_{\mathcal{H}}=0$, since this allows for $g_{w^{*}}(v_{1}^{*})=1-\alpha/2$ and $g_{w^{*}}(-v_{1}^{*})=g_{w^{*}}(v_{-1}^{*})=\alpha/2$ (here $v_{1}^{*}=-v_{-1}^{*}$ since $\pi_{X}$ has mean zero).
-Now we claim that any $w^{\prime}$ with $\norm{w^{\prime}_{\mathcal{H}}}>0$ has strictly greater loss than $w^{*}$.
-To show this, we first observe that $\ell_{\mathrm{LS},\alpha}$ is also convex as a function of $w^{\top}x$ (this is easily check from Equations (A.3) and (A.4)).
-Additionally, we can restrict our attention to $w^{\prime}$ satisfying $\norm{w^{\prime}}\leq M$ for some $M$, since $\ell_{\mathrm{LS},\alpha}(w^{\prime})\to\infty$ as $\norm{w^{\prime}}\to\infty$ for $\alpha>0$.
-Therefore, we can consider $w^{\top}x$ to be integrable and apply the conditional form of Jensen’s inequality to $\ell_{\mathrm{LS},\alpha}(w)$:
-$(\displaystyle\ell_{\mathrm{LS},\alpha}(w)\geq\sum_{y\in\{-1,1\}}-\pi_{Y}(y)\bigg{(}$$\bigg{(}1-\frac{\alpha}{2}\bigg{)}\log\sigma\big{(}\mathbb{E}[yw^{\top}x\mid Y=y]\big{)}$
-$+\frac{\alpha}{2}\log\sigma\big{(}-\mathbb{E}[yw^{\top}x\mid Y=y]\big{)}\bigg{)}$ (A.5)
-The RHS of Equation (A.5) is minimized by our choice of $w^{*}$, and in fact we have equality between $\ell_{\mathrm{LS},\alpha}(w^{*})$ and the RHS.
-Furthermore, since we are dealing with a non-affine function of $w^{\top}x$, we can only achieve equality if $w^{\top}x$ is almost surely constant conditioned on $Y=y$.
-However, by assumption two in Definition 2.1, the conditional variance of $(w_{\mathcal{H}}^{\prime})^{\top}x_{\mathcal{H}}$ is non-zero, and therefore $\ell_{\mathrm{LS},\alpha}(w^{\prime})>\ell_{\mathrm{LS},\alpha}(w^{*})$.
+We follow the outline in the proof sketch.
+We first observe that there exists $M$ such that it suffices to consider $\norm{w}\leq M$; this is due to the fact that for $\alpha>0$, we have $\lim_{\norm{w}\to\infty}\absolutevalue{\ell_{\mathrm{LS},\alpha}(w)}=\infty$.
+With this in mind, let us use $Z$ to denote $w^{T}YX$ and use $\ell_{\mathrm{LS},\alpha}(Z)$ to denote the loss in terms of this quantity.
+Then it follows that
+$\partialderivative{\ell_{\mathrm{LS},\alpha}(Z)}{Z}$ $=\mathbb{E}\left[\sigma(Z)-1+\alpha/2\right],$ (A.3)
+[…]
+where in both cases we applied the dominated convergence theorem, which is justified because $\ell_{\mathrm{LS},\alpha}$ is smooth in $Z$ with bounded derivatives.
+Now since $\norm{w}\leq M$ and the support of $X$ is compact, there exist $\gamma_{1}$ and $\gamma_{2}$ such that $2\partialderivative[2]{\ell_{\mathrm{LS},\alpha}(Z)}{Z}\in(\gamma_{1},\gamma_{2})$, which implies that $\ell_{\mathrm{LS},\alpha}$ is strongly convex in $Z$ and satisfies the conditions of Lemma A.1.
+By Jensen’s inequality, we then have that:
+$\ell_{\mathrm{LS},\alpha}(Z)\geq\left(\frac{\alpha}{2}-1\right)\log\sigma(\mathbb{E}[Z])-\frac{\alpha}{2}\log\sigma(-\mathbb{E}[Z]).$ (A.5)
+Since $\mathbb{E}[YX_{\mathcal{L}}]\neq 0$ by Assumption 3.5, we can choose $w$ such that $w_{\mathcal{H}}=0$ and $\mathbb{E}[Z]=\sigma^{-1}(1-\alpha/2)$, which minimizes the RHS of (A.5).
+Let us use $\mathrm{OPT}$ to denote this minimum.
+Then by Lemma A.1, we have for any $w$ chosen as described:
+$\ell_{\mathrm{LS},\alpha}(w)-\mathrm{OPT}\leq\frac{\gamma_{2}}{2}w^{T}\Sigma_{YX,\mathcal{L}}w.$ (A.6)
+On the other hand, if $w^{\prime}$ is another solution satisfying $\norm{w^{\prime}_{\mathcal{H}}}>\epsilon$, then Lemma A.1 gives
+$\ell_{\mathrm{LS},\alpha}(w^{\prime})-\mathrm{OPT}\geq\frac{\gamma_{1}\rho\epsilon^{2}}{2},$ (A.7)
+from which it is clear that for appropriate $\epsilon$ we cannot have $w^{\prime}$ be a stationary point ($\rho$ above is the same as in Assumption 3.5).
+That we can take $\epsilon\to 0$ as $\norm{\Sigma_{YX,\mathcal{L}}}\to 0$ follows from (A.6).
 ∎
-A.3 Proof of Theorem 3.4
-See 3.4
+See 3.7
 Proof.
-We largely follow the same structure as the proof of Theorem 3.3.
-As before, we compute $\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ and $\gradient_{w}^{2}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ (once again by applying dominated convergence):
-$\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\mathbb{E}_{(x_{1},y_{1}),(x_{2},y_{2})\sim\pi}\left[\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\left(\lambda(g_{w}(y_{1}z_{\lambda})-1)y_{1}+(1-\lambda)(g_{w}(y_{2}z_{\lambda})-1)y_{2}\right)z_{\lambda}\right]\right]$ (A.6)
-$\gradient_{w}^{2}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\mathbb{E}_{(x_{1},y_{1}),(x_{2},y_{2})\sim\pi}\bigg{[}\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\bigg{[}\bigg{(}\lambda(1-g_{w}(y_{1}z_{\lambda}))g_{w}(y_{1}z_{\lambda})$
-$\quad\quad\quad\quad\quad\quad+(1-\lambda)(1-g_{w}(y_{2}z_{\lambda}))g_{w}(y_{2}z_{\lambda})\bigg{)}z_{\lambda}z_{\lambda}^{\top}\bigg{]}\bigg{]}$ (A.7)
-It follows from Equation (A.7) that the Hessian is positive semidefinite (since $\lambda\in[0,1]$), so $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ is also convex.
-Now it also follows from the above calculations that, after conditioning on $\lambda,y_{1},\text{ and }y_{2}$, the function $h_{\lambda}$ from Equation (2.8) (redefined below to indicate that $y_{1},y_{2}$ are also fixed) is convex in $w^{\top}z_{\lambda}$:
-$h_{\lambda,y_{1},y_{2}}(w,z_{\lambda})=-\left(\lambda\log g_{w}(y_{1}z_{\lambda})+(1-\lambda)\log g_{w}(y_{2}z_{\lambda})\right)$ (A.8)
-For convenience, we will use $z_{\lambda}\mid y_{1},y_{2}$ to indicate the conditional distribution of $\lambda x_{1}+(1-\lambda)x_{2}$ given $\lambda,y_{1},y_{2}$ as in Equation (A.8).
-For the moment we suppose that $w^{\top}z_{\lambda}$ is integrable.
-Then by Fubini’s theorem, conditioning, and Jensen’s inequality we obtain:
-$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\mathbb{E}_{y_{1},y_{2}\sim\pi_{Y}}\left[\mathbb{E}_{z_{\lambda}\mid y_{1},y_{2}}\left[h_{\lambda,y_{1},y_{2}}(w,z_{\lambda})\right]\right]\right]$
-$\geq\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\mathbb{E}_{y_{1},y_{2}\sim\pi_{Y}}\left[-\left(\lambda\log\sigma(y_{1}\mathbb{E}[w^{\top}z_{\lambda}])+(1-\lambda)\log\sigma(y_{2}\mathbb{E}[w^{\top}z_{\lambda}])\right)\right]\right]$ (A.9)
-Equality is only obtained in Equation (A.9) if and only if $w^{\top}z_{\lambda}\mid y_{1},y_{2}$ is almost surely constant, which only occurs if $\norm{w_{\mathcal{H}}}=0$ as before.
-It remains to show in this case that we can obtain a stationary point with $\norm{w_{\mathcal{H}}}=0$, as then that will imply that any choice of $w$ with $\norm{w_{\mathcal{H}}}>0$ will have strictly worse loss.
-To do so, let us consider $w$ with $\norm{w_{\mathcal{H}}}=0$ and $\left\langle w,v_{1}^{*}\right\rangle=u=-\left\langle w,v_{-1}^{*}\right\rangle$.
-Let $\mathbb{P}_{\lambda}$ denote the probability measure associated with $\mathcal{D}_{\lambda}$.
-We then define the following two functions, which will correspond to the terms in $\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ that result from mixing $y_{1}=y_{2}$ and $y_{1}\neq y_{2}$.
-$f_{1}(u)$ $=\sigma(u)-1$ (A.10)
-$f_{2}(u,\lambda)$ $=\sigma((2\lambda-1)u)-\lambda$ (A.11)
-Now, recalling that $\mu_{y}=\mathbb{E}[X\mid Y=y]$ and that $\mu_{1}=-\mu_{-1}$ since $\pi_{Y}(1)=\pi_{Y}(-1)$ and $\mathbb{E}_{x\sim\pi_{X}}[x]=0$, we get:
-$\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\mathbb{E}_{y_{1},y_{2}\sim\pi_{Y}}\left[\mathbb{E}_{z_{\lambda}\mid y_{1},y_{2}}\left[\left(\lambda(g_{w}(y_{1}z_{\lambda})-1)y_{1}+(1-\lambda)(g_{w}(y_{2}z_{\lambda})-1)y_{2}\right)z_{\lambda}\right]\right]\right]$
-$=\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\frac{1}{2}\mathbb{E}_{z_{\lambda}\mid y_{1}=1,y_{2}=1}\left[f_{1}(u)z_{\lambda}\right]+\frac{1}{2}\mathbb{E}_{z_{\lambda}\mid y_{1}=1,y_{2}=-1}\left[f_{2}(u,\lambda)z_{\lambda}\right]\right]$
+Let us first outline the overall steps of the proof, and the differences with the label smoothing case.
+  1. 1.
+    We first show that the loss conditioned on $\lambda$ is strongly convex in $w^{\top}Z_{\lambda}$.
+The conditioning on $\lambda$ here is necessary because $\lambda$ is a random variable, unlike $\alpha$ in the label smoothing case.
+The overall goal here is to use the same argument as for label smoothing, i.e.
+show that we can achieve the optimal lower bound in terms of $\mathbb{E}[w^{\top}Z_{\lambda}]$ using only $w_{\mathcal{L}}$ and letting $\norm{\Sigma_{YX,\mathcal{L}}}\to 0$.
+  2. 2.
+    We cannot explicitly minimize the conditional loss like we did with label smoothing, since it is not possible with a fixed choice of $w$ to achieve $\sigma^{-1}(\mathbb{E}[w^{\top}Z_{\lambda}])=\lambda$ for every $\lambda$ simultaneously.
+Instead, we will show that a stationary point of the conditional loss exists that uses only the dimensions of $w_{\mathcal{L}}$.
+  3. 3.
+    Having shown the above, we can just reuse the same argument as before with Lemma A.1 to prove the desired result.
+Let $\ell_{\mathrm{MIX},\lambda}$ denote $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$ with a fixed choice of $\lambda$ (i.e.
+after conditioning on $\lambda$), and let $R=w^{\top}Z_{\lambda}$.
+Then we can compute:
+$\partialderivative{\ell_{\mathrm{MIX},\lambda}(R)}{R}$ $=\mathbb{E}\big{[}\lambda(\sigma(Y_{1}R)-1)Y_{1}+(1-\lambda)(\sigma(Y_{2}R)-1)Y_{2}\big{]},$ (A.8)
 […]
-Where in the penultimate step we used the fact that the terms that result from conditioning with $y_{1}=y_{2}$ are the same after substituting $\mu_{-1}=-\mu_{1}$, and likewise for the mixed term.
-Now we consider the limiting behavior of Equation (A.12) as $u\to\pm\infty$.
-We first focus on the integral term in Equation (A.12), which we can analyze via dominated convergence:
+where again we applied dominated convergence to the expectation with respect to $\pi$.
+Strong convexity follows from the same consideration as label smoothing; namely, we can consider $\norm{w}\leq M$ as $\ell_{\mathrm{MIX},\lambda}(R)\to\infty$ as $\norm{w}\to\infty$ so long as $\lambda$ is not 0 or 1 and $\pi_{Y}(y)>0$, and this consequently implies (A.9) is lower bounded by some positive real number.
+Now by conditional Jensen’s inequality, we obtain the following lower bound for $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$:
+$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)\geq\mathbb{E}_{\lambda}\left[\mathbb{E}_{Y_{1},Y_{2}}\left[\ell_{\mathrm{MIX},\lambda}(\mathbb{E}\left[R\mid\lambda,Y_{1},Y_{2}\right])\right]\right].$ (A.10)
+We now show that it is possible to minimize this lower bound while taking $w_{\mathcal{H}}=0$.
+This is more difficult than it was in the label smoothing case, because it is no longer obvious that $\mathbb{E}[YX_{\mathcal{L}}]\neq 0$ is sufficient for minimizing the RHS of (A.10) due to the expectation with respect to $\lambda$.
+However, we can show the existence of a stationary point with $w_{\mathcal{H}}=0$, even though we cannot provide an explicit construction.
+The idea is to consider the limiting behavior of (A.10) as we take the values of $w_{\mathcal{L}}$ to $\infty$- ∞ and $\infty$.
+Note that we can take this limit into the expectation with respect to $\lambda$ by dominated convergence again.
+Let us consider the gradient with respect to $w$ of the RHS of (A.10).
+To make notation manageable, we will use $a_{1}=\mathbb{E}[X\mid Y=1],a_{2}=\mathbb{E}[X\mid Y=-1]$, and $a_{3}=\mathbb{E}[Z_{\lambda}\mid\lambda,Y_{1}=1,Y_{2}=-1]$.
+We can then explicitly write out the gradient as the expectation with respect to $\lambda$ of the sum of the following three terms (considering the different cases for $Y_{1},Y_{2}$):
+$\gradient_{w}\mathbb{E}_{Y_{1},Y_{2}}\left[\ell_{\mathrm{MIX},\lambda}(\mathbb{E}\left[R\mid\lambda,Y_{1},Y_{2}\right])\right]$ $=\pi_{Y}(1)^{2}\left(\sigma(w^{\top}a_{1})-1\right)a_{1}$
+$\quad-\pi_{Y}(-1)^{2}\left(\sigma(-w^{\top}a_{2})-1\right)a_{2}$
+$\quad+2\pi_{Y}(1)\pi_{Y}(-1)\bigg{(}\lambda\big{(}\sigma(w^{\top}a_{3})-1\big{)}$
+$\quad\quad\quad-(1-\lambda)\big{(}\sigma(-w^{\top}a_{3})-1\big{)}\bigg{)}a_{3}.$ (A.11)
+The first two lines above are obtained from the fact that we can combine terms when $Y_{1}=Y_{2}$, and the last line is by symmetry.
+Now we recall that by assumption $\mathbb{E}[YX_{\mathcal{L}}]\neq 0$ and $\mathbb{E}[X]=0$.
+Thus, WLOG, we can assume that $\mathbb{E}[X_{\mathcal{L}}\mid Y=1]_{i}>0$ and $\mathbb{E}[X_{\mathcal{L}}\mid Y=-1]_{i}<0$ for each index $i$.
+With this in mind, we consider first the case of what happens when the entries $w_{\mathcal{L}}\to\infty$.
+Since $w^{\top}a_{1}>0$ and $w^{\top}a_{2}<0$, the first two terms in (A.11) vanish (independent of $\lambda$).
+On the other hand, for the third term, there are two cases to consider.
+Depending on $\lambda$, we have that the entries of $a_{3}$ are either strictly negative or strictly positive, with the exceptional case of $\lambda=\pi_{Y}(1)$ in which $a_{3}=0$.
+If the entries of $a_{3}$ are strictly negative, then $\big{(}\sigma(-w^{\top}a_{3})-1\big{)}\to 0$ and the coefficient becomes negative, so the third term is positive.
+Similarly, if $a_{3}$ is strictly positive, the coefficient is positive and the third term is still positive.
+Thus, as $w_{\mathcal{L}}\to\infty$ every entry of (A.11) is positive.
 […]
-$\quad+\int_{0.5}^{1}\big{(}\sigma((2\lambda-1)u)-\lambda\big{)}(2\lambda-1)\ d\mathbb{P}_{\lambda}(\lambda)\bigg{)}$
-[…]
-$\geq 0$ (A.13)
-And similarly, the limit as $u\to-\infty$ is bounded above by 0. Thus, from Equation (A.12) we get that:
-$\lim_{u\to\infty}\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\beta_{1}\mu_{1}\quad\text{for some}\quad\beta_{1}\geq 0$ (A.14)
-$\lim_{u\to-\infty}\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\beta_{2}\mu_{1}\quad\text{for some}\quad\beta_{2}\leq 0$ (A.15)
-Since Equation (A.12) is continuous in $u$, it then follows that there exists $u$ such that $\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)=0$, which implies that we can make our choice of $w$ a stationary point.
-However, there is one subtlety here - the fact that any other $w^{\prime}$ with $\norm{w^{\prime}_{\mathcal{H}}}>0$ has worse loss hinges on the use of Jensen’s inequality above, which requires that we constrain ourselves to the case where $w^{\top}z_{\lambda}$ is integrable.
-[…]
-For this case, if $\norm{w^{\prime}_{\mathcal{H}}}>0$, then $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w^{\prime})\to\infty$ as $\norm{w^{\prime}}\to\infty$, since we incur infinite loss on the mixed terms $y_{1}\neq y_{2}$.
-As a result, we can conclude that every minimizer of $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$ (for $\mathcal{D}_{\lambda}$ satisfying the conditions of the theorem) has zero norm in the high variance dimensions, including those minimizers that occur as the norm of the weights tend to infinity.
+Now we have basically arrived at the same stage as the end of the label smoothing proof.
+By taking $\norm{\Sigma_{YX,\mathcal{L}}}\to 0$ we can get arbitrary concentration around this optimal $R$, and by the same logic as the label smoothing proof the result follows.
 ∎
-Appendix B Low Variance Experiments
-(a) Weight Decay
-(b) Label Smoothing
-(c) Mixup
-Figure 6: Final weight norm of logistic regression model in the high variance dimensions ($\norm{w_{\mathcal{H}}}$) for various hyperparameter settings on the synthetic data distribution of Section 4.1 modified so that the dimensions in $\mathcal{L}$ have non-zero variance.
-(a) Weight Decay
-(b) Label Smoothing
-(c) Mixup
-Figure 7: Logistic regression final test errors for various hyperparameter settings on the binary classification versions of CIFAR-10 and CIFAR-100 from Section 4.2, but with the training data modified to have non-zero variance in the identifying dimension.
-Figures 6 and 7 correspond to non-zero variance versions of the experiments of Sections 4.1 and 4.2.
-In Figure 6, we consider synthetic data as per Definition 4.1 except for that $x_{i}=y(\gamma+\xi)$ where $\xi\sim\mathrm{Uniform}([0,\gamma/10])$ for $i\leq\lfloor d/2\rfloor$ (i.e.
-the $\mathcal{L}$ dimensions now have non-zero variance).
-Similarly, Figure 7 considers the same setting as Section 4.2 but with the first dimension of the input replaced with $y(\gamma+\xi)$ instead of $y\gamma$.
-As can be seen from the results, the introduction of a small amount of variance to these settings leaves model behavior effectively unchanged.
-We also considered a similar low variance setup under the setting of Section 4.3.
-To do so, we modify the zero variance concatenated features with added mean-zero Gaussian noise in the training data.
-However, as we noted in the main paper, we found that unless the scaling of the added Gaussian noise was significantly less than $\gamma$ (i.e.
-$\gamma*10^{-3}$), the label smoothing and Mixup overfitting phenomenon was also exhibited by weight decay.
-This is possibly just a consequence of our experimental setup and/or the selected hyperparameters (i.e.
-perhaps this sensitivity is avoidable with some optimal choice of the weight decay hyperparameter), but we do not exhaustively explore this and instead leave a more comprehensive understanding of this phenomenon to follow-up work.
+A.3 Proofs for Section 3.2
+See 3.9
+Proof.
+Let us first define:
+$\ell_{\mathrm{LS},\alpha,y}(g)=-\mathbb{E}\bigg{[}(1-\alpha)\log g^{y}(X)+\frac{\alpha}{k}\sum_{i=1}^{k}\log g^{i}(X)\mid Y=y\bigg{]}.$ (A.12)
+Then we can decompose $\ell_{\mathrm{LS},\alpha}(g)$ as follows:
+$\ell_{\mathrm{LS},\alpha}(g)=\sum_{y=1}^{k}\pi_{Y}(y)\ell_{\mathrm{LS},\alpha,y}(g).$ (A.13)
+Once again, since we can restrict our attention to $g(X)\in[\gamma,1-\gamma]$ for some $\gamma$ (as the loss goes to infinity if $g(X)=1$ on a set of positive $\pi_{X}$-measure with $\alpha>0$), it is easy to verify that (A.12) is strongly convex in $g(X)$.
+The desired result then follows by applying Lemma A.1 with the regular conditional distribution $g(X)\mid Y=y$ for each term in (A.13).
+∎
+See 3.10
+Proof.
+The proof follows an identical structure to that of Proposition 3.9. In particular, we again define the following conditional loss:
+$\ell_{\mathrm{MIX},\lambda,y_{1},y_{2}}=-\mathbb{E}[\lambda\log g^{y_{1}}(Z_{\lambda})+(1-\lambda)\log g^{y_{2}}(Z_{\lambda})\mid y_{1},y_{2},\lambda].$ (A.14)
+And we can then decompose $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$ as:
+$\ell_{\mathrm{MIX},\lambda}(g)=\sum_{y_{1}=1}^{k}\sum_{y_{2}=1}^{k}\pi_{Y}(y_{1})\pi_{Y}(y_{2})\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}[\ell_{\mathrm{MIX},\lambda,y_{1},y_{2}}(g)].$ (A.15)
+Since $\mathcal{D}_{\lambda}$ is not a point mass on 00 or $1$, we can restrict ourselves to $g(X)\in[\gamma,1-\gamma]$ for some $\gamma$ as before, and strong convexity in $g(X)$ of the conditional loss (A.14) again follows.
+We then apply Lemma A.1 to obtain the result.
+∎

### v3

@@
 We then use our techniques to show that even for nonlinear models and general data distributions, the label smoothing and Mixup losses are lower bounded by a function of the model output variance.
-An important consequence of our results is negative: label smoothing and Mixup can be less robust to spurious correlations in the data.
-We verify that our theory reflects practice via experiments on image classification benchmarks modified to have spurious correlations.
+Lastly, we demonstrate empirically that this aspect of label smoothing and Mixup can be a positive and a negative.
+On the one hand, we show that the strong performance of label smoothing and Mixup on image classification benchmarks is correlated with learning low variance hidden representations.
+On the other hand, we show that Mixup and label smoothing can be more susceptible to low variance spurious correlations in the training data.
 1 Introduction
@@
 In this paper, we aim to provide such a perspective by extending the ideas from previous feature learning analyses to the aspect shared in common between both Mixup and label smoothing: label augmentation.
-Our analysis allows us to identify a clear failure mode shared by both Mixup and label smoothing - that they can ignore higher variance features - which we also verify empirically.
+Our analysis shows that both Mixup and label smoothing hone in on low variance features in the data, and we demonstrate empirically that this phenomenon occurs in both synthetic settings with spuriously correlated features as well as in the training of standard deep learning models for image classification benchmarks.
 1.1 Main Contributions
-Our main contributions are three-fold:
-  1. 1.
-    We introduce a simple class of data distributions in which there exists a low variance feature and a high variance feature, and show that optimizing the label smoothing and Mixup losses for a linear model for such distributions leads to learning only the low variance feature.
-On the other hand, if the high variance feature satisfies a greater separability condition relative to the low variance feature, optimizing the empirical risk with weight decay successfully learns both features in the data.
-Our label smoothing and Mixup results are stated under general conditions on both the data and the respective loss hyperparameters; our Mixup results apply to any mixing distribution that does not coincide with ERM, which greatly generalizes the settings considered by prior work on Mixup feature learning (Chidambaram et al., 2023; Zou et al., 2023).
-  2. 2.
-    We show that our techniques extend to the analysis of general model classes, and prove lower bounds for the label smoothing and Mixup losses in terms of the model output variance.
-These results suggest that optimizing these losses necessarily decreases model output variance, which is not necessarily the case for training with weight decay.
-  3. 3.
-    We empirically verify our theoretical insights by considering image classification benchmarks modified to (1) have adversarial spurious correlations in the training data and (2) more natural spurious correlations in the form of class-correlated backgrounds.
-In both cases, we observe that label smoothing and Mixup fail to generalize for the full range of hyperparameters used in practice, while ERM with weight decay does generalize.
+The main message of our paper can be summarized as:
+In data distributions where there are low variance and high variance latent features, both label smoothing and Mixup incentivize learning models that correlate more strongly with the low variance features.
+We prove this message concretely in Section 3.1, where we characterize linear optimizers of the label smoothing and Mixup losses on data in which some dimensions have lower variance than others.
+In Section 3.2, we also prove weaker analogues of our linear results in the context of general models and multi-class distributions; namely, we show that optimizing the label smoothing and Mixup losses requires decreasing model output variance in a way that is not necessarily true for empirical risk minimization (ERM) combined with weight decay.
+We verify our theory empirically in Section 4, where we show that label smoothing and Mixup do indeed learn lower variance hidden representations (corroborating ideas from the literature on neural collapse (Papyan et al., 2020)) and that this lower variance property correlates with better generalization performance.
+We hypothesize that, for standard benchmarks, there exist low variance latent features that generalize well as opposed to high-variance, noisy features, and we expect that directly investigating such features would be a fruitful line of future work.
+However, we also point out that our key message does not directly imply better generalization performance – we show via synthetic experiments in Section 4.2 that it is possible for the low variance features to actually be spuriously correlated with the targets (e.g.
+fixed backgrounds).
 1.2 Related Work
@@
 (2015), label smoothing continues to be used to train SOTA vision (Wortsman et al., 2022; Liu et al., 2022), translation (Vaswani et al., 2017; Team et al., 2022), and multi-modal models (Yu et al., 2022).
-Attempts to understand when and why label smoothing is effective can be traced back to the influential empirical analyses of Pereyra et al.
+Attempts to understand when and why label smoothing is effective can be traced back to Pereyra et al.
 (2017) and Müller et al.
-(2020), which respectively relate label smoothing to entropy regularization and show that it can lead to more closely clustered learned representations, with the latter work further showing that these clustered representations can be responsible for decreased distillation performance.
+(2020), which respectively relate label smoothing to entropy regularization and show that it can lead to more closely clustered learned representations.
+Müller et al.
+(2020) also show that label smoothing can improve calibration but hurt distillation performance; Zhu et al.
+(2023); Xia et al.
+(2024) further show that these improvements in calibration do not translate to better selective classification.
 On the theoretical side, Lukasik et al.
@@
 (2020) provide an alternative theoretical perspective, showing that label smoothing can improve convergence of stochastic gradient descent.
-However, to the best of our knowledge, there are no preexisting theoretical works attempting to understand how label smoothing affects feature learning.
 Mixup.
@@
 (2021) show that the regularization terms introduced by the Mixup loss can also have a negative impact due to the fact that Mixup-augmented points may coincide with existing training data points, leading to models that fail to minimize the original risk.
-Mixup has also been studied from the perspective of calibration (Thulasidasan et al., 2019), with theoretical results (Zhang et al., 2021; Chidambaram & Ge, 2023) showing that Mixup training can prevent miscalibration issues arising from ERM (even when post-training calibration methods are used with ERM).
+Mixup has also been studied from the perspective of calibration (Thulasidasan et al., 2019), with theoretical results (Zhang et al., 2021; Chidambaram & Ge, 2023) showing that Mixup training can prevent miscalibration issues arising from ERM.
 Recently, Oh & Yun (2023) studied Mixup in a similar context to our work (binary classification with linear models), and showed that Mixup can significantly improve the sample complexity required to learn the optimal classifier when compared to ERM.
@@
 (2023) both showing that Mixup training can learn multiple features in a generative, multi-view (Allen-Zhu & Li, 2021) data model despite ERM failing to do so.
-We provide a more detailed comparison to these works when we discuss our main results in Section 3.
+Neural Collapse.
+Papyan et al.
+(2020) showed that the last layer activations of trained neural networks collapse to their class means, in such a way that they are maximally separable.
+Kornblith et al.
+(2021); Zhou et al.
+(2022); Guo et al.
+(2024) all investigate the interplay between label smoothing and this effect, with Kornblith et al.
+(2021) showing that label smoothing can lead to more separable last layer representations (although worse linear transfer performance) and Zhou et al.
+(2022); Guo et al.
+(2024) showing that it can also lead to faster convergence to these collapsed representations.
+We corroborate these results with our experiments in Section 4.1, where we show that both Mixup and label smoothing lead to much lower total variance in last layer activations compared to standard cross-entropy.
 General Data Augmentation.
@@
 After which we can define the Mixup cross-entropy as:
-$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(g)=\mathbb{E}_{(X_{1:2},Y_{1:2})\sim\pi,\lambda\sim\mathcal{D}_{\lambda}}\left[-h(\lambda,g,X_{1:2},Y_{1:2})\right].$ (2.8)
+$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(g)=\mathbb{E}_{(X_{1:2},Y_{1:2})\sim\pi\otimes\pi,\lambda\sim\mathcal{D}_{\lambda}}\left[-h(\lambda,g,X_{1:2},Y_{1:2})\right].$ (2.8)
 The corresponding binary version $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ of (2.8) is identical except for redefining $h$ to be:
@@
 We begin first with the binary setting, in which we will consider a data model where a subset of the input dimensions correspond to a low variance feature and the complementary dimensions correspond to a high variance feature that is more separable.
-This setup is intended to be a simplistic model of spurious correlations that appear in image data; one could think of the low variance feature as a fixed background while the high variance feature corresponds to the actual object category of interest.
-The goal in such cases is to ensure that we do not only learn the low variance feature, so that generalization is possible even under a shift of this feature (for example, recognizing the same object on different backgrounds).
-We will show, however, that doing any kind of label augmentation (label smoothing or Mixup) in this low variance/high variance feature setup will lead to a model that has only learned the low variance feature, whereas minimizing the binary cross-entropy with non-zero weight decay requires learning the high variance feature due to its greater separability.
-The overarching data model for this subsection is as follows.
+We emphasize that our notion of “feature” here does not correspond to explicit feature vectors that are fixed per class like in prior work, we simply designate subsets of the dimensions as features for simplicity.
+This data model can be interpreted in multiple ways: depending on the context, we may wish to learn either both of the features present in the data or simply hone in on the low variance feature (for example, identifying a stop sign with many different backgrounds).
+Although this setting seems simplistic – one may object that the variance differences in the input can be handled with suitable normalization/whitening – the insights from our data model can be applied to learned features (i.e.
+intermediate representations in a deep learning model) where such modifications are less straightforward, as we demonstrate in the experiments of Section 4.
+Our main results in this section show that doing any kind of label augmentation (label smoothing or Mixup) in our low variance/high variance feature setup will lead to a model that has only learned the low variance feature, whereas minimizing the binary cross-entropy with non-zero weight decay requires learning the high variance feature due to its greater separability.
 Definition 3.1.
@@
 For convenience, for a vector $v\in\mathbb{R}^{d}$, we will use $v\in\mathcal{L}$ to mean that only the $\mathcal{L}$ dimensions of $v$ are non-zero.
-Definition 3.1 does not actually make any strong assumptions on $\pi$ yet; this is because the weight decay and label augmentation results have different, non-conflicting assumptions that we codify separately.
-For the former, we need the data to be more separable in the $\mathcal{H}$ dimensions than in the $\mathcal{L}$ dimensions, in the following sense.
+Definition 3.1 only treats $\mathcal{L}$ and $\mathcal{H}$ as placeholders; this is because our weight decay result is insensitive to variance assumptions and only depends on differences in the separability of the dimensions $\mathcal{L}$ and $\mathcal{H}$, as we indicate below.
 Assumption 3.2.
@@
 Assumption 3.2 is strong, but it actually does not imply linear separability, only that the $\mathcal{H}$ dimensions are in a sense better than the $\mathcal{L}$ dimensions for classification.
-For a simple example to provide intuition for the assumption, one can consider a 2-D distribution in which the $x$ coordinates for both classes are concentrated close to 0, but the $y$ coordinates are either large negative or large positive values depending on the class.
+We explore a simple 2-D distribution illustrating Assumption 3.2 in Appendix C.1.
 Our first result shows that for $\ell_{\beta}(w)$ as defined in (2.2), the minimizer $w^{*}$ has a large correlation with the dimensions in $\mathcal{H}$.
@@
 If $w^{*}$ is the max-margin solution to $\ell_{0}(w)$, then the result of Theorem 3.3 still holds.
-On the other hand, as alluded to earlier, this phenomenon does not occur when considering the label smoothing and Mixup losses $\ell_{\mathrm{LS},\alpha}$ and $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$.
-We will show that for both label smoothing and Mixup, the optimal solutions have arbitrarily small correlation with the dimensions in $\mathcal{H}$ as the distribution of $YX$ concentrates.
+On the other hand, we will now show that once we introduce variance assumptions on $\mathcal{L}$ and $\mathcal{H}$, this phenomenon does not occur when minimizing the label smoothing and Mixup losses $\ell_{\mathrm{LS},\alpha}$ and $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$.
+In both cases, the optimal solutions have arbitrarily small correlation with the dimensions in $\mathcal{H}$ as the distribution of $YX$ concentrates.
+[^1] In order for the variance of $YX$ to go to 00, we need the classes to be balanced; however, we can change our assumptions to be in terms of the conditional variance of $X$ to remove this requirement.
 For these results, we require the following assumptions (but no longer need Assumption 3.2).
 Assumption 3.5.
-We assume that $\mathbb{E}[YX_{\mathcal{L}}]\neq 0$ and $\Sigma_{YX,\mathcal{H}}\succ\rho\mathrm{Id}$ for some $\rho>0$.
+We assume that $\mathbb{E}[YX_{\mathcal{L}}]\neq 0$ and $\Sigma_{YX,\mathcal{H}}\succ\rho\mathrm{I}_{d}$ for some $\rho>0$.
 Here $\Sigma_{YX,\mathcal{H}}$ denotes the covariance matrix of $\Sigma_{YX}$ restricted to those rows and columns in $\mathcal{H}$.
@@
 Observe that we have made no direct separability assumptions on the data, although it is true that as $\norm{\Sigma_{YX,\mathcal{L}}}\to 0$ the class-conditional supports are guaranteed to be linearly separable in $\mathcal{L}$.
-We now state both results.
+This means that we can consider settings where both Assumptions 3.2 and 3.5 are true, and in such settings there will be a clear separation between weight decay and label smoothing/Mixup.
+This kind of separation in the learned decision boundaries is visualized in Appendix C.1, which provides intuition for the next two results.
 Theorem 3.6.
@@
 Proof Sketch.
-The proof follows the same outline as the label smoothing result, but there are now additional difficulties due to mixing inputs.
-In particular, we need to condition on $\lambda$ and then show strong convexity of the conditional loss in terms of $w^{\top}Z_{\lambda}$.
-We can then get a lower bound on the overall loss as an expectation with respect to $\lambda$ of this conditional loss.
-However, it is no longer immediate that there exists a stationary point $w\in\mathcal{L}$ that minimizes this lower bound.
-We prove the existence of such a stationary point by considering limiting behavior of the gradient of the loss as $w_{\mathcal{L}}$ tends to $\infty$ or $\infty$- ∞ in each component.
-After establishing a stationary point, we can show as with the label smoothing result that as $\norm{\Sigma_{YX,\mathcal{L}}}\to 0$ we can get a solution with $w\in\mathcal{L}$ that achieves the lower bound while solutions with $w_{\mathcal{H}}$ non-zero are bounded away from the lower bound.
+Mixup differs from label smoothing in that we need to condition on $\lambda$ and then show strong convexity of the conditional loss in terms of $w^{\top}Z_{\lambda}$.
+We can then get a lower bound similar to the label smoothing case, but it is no longer immediate that there exists a stationary point $w\in\mathcal{L}$ that minimizes this lower bound.
+We prove the existence of such a stationary point by considering limiting behavior of the gradient of the loss as $w_{\mathcal{L}}$ tends to $\infty$ or $\infty$- ∞ in each component, after which the rest of the proof follows the label smoothing proof.
 3.2 General Multi-Class Classification
@@
 For $\alpha>0$ and any $g:\mathbb{R}^{d}\to\Delta^{k-1}$, letting $\mathrm{OPT}_{\mathrm{LS},\alpha}$ denote the minimum of $\ell_{\mathrm{LS},\alpha}$, we have for a universal constant $C>0$:
-$\ell_{\mathrm{LS},\alpha}(g)\geq\mathrm{OPT}_{\mathrm{LS},\alpha}+C\sum_{y=1}^{k}\pi_{Y}(y)\mathrm{Var}\left(g(X)\mid Y=y\right).$ (3.1)
+$\ell_{\mathrm{LS},\alpha}(g)\geq\mathrm{OPT}_{\mathrm{LS},\alpha}+C\sum_{y=1}^{k}\pi_{Y}(y)\mathrm{Var}\left(\mathbb{E}[g(X)\mid Y=y]\right).$ (3.1)
 Proposition 3.10.
 For any $\mathcal{D}_{\lambda}$ that is not a point mass on 00 or $1$ and any $g:\mathbb{R}^{d}\to\Delta^{k-1}$, letting $\mathrm{OPT}_{\mathrm{MIX},\mathcal{D}_{\lambda}}$ denote the minimum of $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$, we have for a universal constant $C>0$ that:
-$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(g)\geq\mathrm{OPT}_{\mathrm{MIX},\mathcal{D}_{\lambda}}+C\sum_{y_{1}=1}^{k}\sum_{y_{2}=1}^{k}\pi_{Y}(y_{1})\pi_{Y}(y_{2})\mathbb{E}_{\lambda}\left[\mathrm{Var}\left(g(Z_{\lambda})\mid y_{1},y_{2},\lambda\right)\right].$ (3.2)
+$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(g)\geq\mathrm{OPT}_{\mathrm{MIX},\mathcal{D}_{\lambda}}+C\sum_{y_{1}=1}^{k}\sum_{y_{2}=1}^{k}\pi_{Y}(y_{1})\pi_{Y}(y_{2})\mathbb{E}_{\lambda}\left[\mathrm{Var}\left(\mathbb{E}[g(Z_{\lambda})\mid y_{1},y_{2},\lambda]\right)\right].$ (3.2)
 The proofs for both Propositions 3.9 and 3.10 follow the same structure; we show that after appropriate conditioning, both losses can be broken up into a sum of strongly convex conditional losses.
@@
 (2023), which were that Mixup can learn both a “common” feature and a “rare” feature in the data; in their setup the common/rare features are fixed (zero variance) per class and concatenated with high variance noise.
+We provide a more detailed review of the settings of Chidambaram et al.
+(2023) and Zou et al.
+(2023) in Appendix B.
 4 Experiments
-We now verify the predictions of our theoretical results in Section 3 across both binary and multi-class experimental setups.
-All experiments in this section were conducted on a single P100 GPU using PyTorch (Paszke et al., 2019) for model implementation.
+We now address the practical ramifications of our theoretical results.
+In Section 4.1, we show that the intermediate representations learned by deep learning models trained with label smoothing and Mixup do indeed exhibit significantly lower variance when compared to those learned using just weight decay, and that these lower variance representations correlate with better performance.
+We also show, however, that this minimum variance feature learning can be a detriment by analyzing spurious correlations in the training data in Section 4.2.
+We also include synthetic experiments directly verifying our theoretical setting in Appendix C.2.
+All experiments in this section were conducted on a single A5000 GPU using PyTorch (Paszke et al., 2019) for model implementation.
 All reported results correspond to means over 5 training runs, and shaded regions in the figures correspond to 1 standard deviation bounds.
-4.1 Spurious Pixels in Binary Classification
+4.1 Learned Low Variance Features in Image Classification
+(a) CIFAR-10 Test Error
+(b) CIFAR-10 Activation Variance
+(c) CIFAR-10 Output Variance
+(d) CIFAR-100 Test Error
+(e) CIFAR-100 Activation Variance
+(f) CIFAR-100 Output Variance
+Figure 1: ResNet-18 final test errors, penultimate layer activation variances, and output probability variances on CIFAR-10 and CIFAR-100.
+Activation variance results are shown starting at epoch 25 as early epochs have larger scale oscillations in the computed variance.
+We first consider image classification on the standard benchmarks of CIFAR-10 and CIFAR-100 (Krizhevsky, 2009) using ResNets (He et al., 2015); we show results for ResNet-18 in this section and relegate further experiments for deeper architectures to Appendix C.3 (they follow the same trends).
+We compare the performance of training using just ERM + weight decay to that of ERM + weight decay combined with label smoothing or Mixup; final test error performance is shown in Figure 1 (a) and (d).
+Due to compute constraints, we focus on known well-performing settings (Zhang et al., 2018; Müller et al., 2020) for weight decay, label smoothing, and Mixup.
+Namely, we take the weight decay parameter to be $5\times 10^{-4}$, the label smoothing $\alpha$ parameter to be $0.1$, and the mixing distribution for Mixup to be $\mathrm{Beta}(1,1)$ (uniform distribution).
+We train all models for 200 epochs using a batch size of 1024, a fixed learning rate of $10^{-3}$, and AdamW for optimization (Loshchilov & Hutter, 2019).
+We preprocess the training data to have zero mean and unit variance along each channel, and also include random crop and flip augmentations as is standard practice for achieving good performance.
+Figure 1 shows that adding in label smoothing and Mixup to the baseline of ERM + weight decay leads to noticeable performance improvements, which corroborates existing results in the literature.
+The novel aspect of our results is shown in Figure 1 (b), (c), (d), and (e), where we track the mean total variance of the penultimate layer activations and output probabilities of each model on the test data over the course of training.
+The variance values are computed by first computing the total variance of the activations/probabilities (i.e.
+the sum of the variance in each dimension) for each class, and then averaging over all classes.
+Final test error and variance values are provided in Table 1 and 2 in Appendix C.3.1.
+Essentially, we measure the average spread of activations and probabilities across classes, which is similar in spirit to what was done by Müller et al.
+(2020), although we directly look at variances across all classes whereas they look at projected clusters of a few specific classes.
+The most telling results are the model activation variances, since it is to an extent expected that per-class variance of model outputs should decrease with improvements in test error (this also corresponds to improved model calibration, which is a known consequence of training with label smoothing and Mixup).
+That being said, we further analyze model output variance in Appendix C.4 and show that label smoothing and Mixup decrease total output variance by a larger extent than what can be explained by changes in the target class prediction variance alone.
+Overall, our results show that adding either label smoothing or Mixup to the baseline of just weight decay leads to significant decreases in the activation variances, adding credence to the idea that both methods learn low variance features.
+We note, however, that our results only establish a correlation between this low variance property and better test performance – it would require a significantly more in-depth empirical study to assess a causal relationship between this kind of feature learning and generalization, which we think would be a fruitful avenue for future work.
+4.2 Low Variance Spurious Correlations
+We now demonstrate that honing in on low variance features can also be harmful to performance via binary classification and multi-class classification tasks in which the training data is modified to have spurious correlations with the target.
+The introduced spurious correlations are much lower magnitude than the rest of the data, and in that sense they intuitively satisfy both Assumptions 3.2 and 3.5.
+4.2.1 Binary Classification With Perturbed Training Data
 (a) Weight Decay
@@
 (c) Mixup
-Figure 1: Logistic regression final test errors for various hyperparameter settings on the binary classification versions of CIFAR-10 and CIFAR-100 from Section 4.1.
-We first consider the case of logistic regression trained on data that has trivial spurious correlations that do not appear in the test data.
-For our binary classification task, we consider reductions of the standard CIFAR-10 and CIFAR-100 (Krizhevsky, 2009) benchmarks to binary classification by fixing two classes from each dataset as the positive and negative classes, and replacing their original labels with the labels $+1$ and $-1$ respectively.
-This leads to 10000 training examples and 2000 test examples in the case of CIFAR-10, and 1000 training examples and 200 test examples in the case of CIFAR-100.
-Our experiments are not sensitive to the choice of binary reduction, and for the experiments in this section we fix the $-1$ class to be class 00 from the original data and the $+1$ class to be class $1$ from the original data (“airplane” and “automobile” for CIFAR-10, “apple” and “aquarium fish” for CIFAR-100).
-We preprocess the training data to have mean zero and variance one along every image channel (and apply the same transform to the test data).
-To introduce spurious correlations into the training data, we modify the data such that the first value in the tensor representation of each training input (corresponding to the “R” channel of the top left pixel in the original image) is replaced by $y\gamma$ with $\gamma=0.1$ (the choice of $\gamma$ here is not important as long as it is small, we verified the results for $\gamma=10^{-5}$ up to $\gamma=0.1$).
-This ensures that the training data is trivially linearly separable in the first dimension of the data, but learning this first dimension requires having larger weight norm due to the scaling by $\gamma$.
+Figure 2: Logistic regression final test errors for various hyperparameter settings on the binary classification versions of CIFAR-10 and CIFAR-100 from Section 4.2.1.
+For our binary classification task, we consider reductions of the CIFAR benchmarks to binary classification by fixing two classes from each dataset as the positive and negative classes, and replacing their original labels with the labels $+1$ and $-1$ respectively.
+Our experiments are not sensitive to the choice of binary reduction, and for the experiments in this section we fix the $-1$ class to be class 00 from the original data and the $+1$ class to be class $1$ from the original data.
+We preprocess the training data to have mean zero and variance one along every image channel, but do not use other augmentations or perform feature extraction using pretrained models.
+We then “adversarially” modify the training data such that the first value in the tensor representation of each training input is replaced by $\gamma y$ with $\gamma=0.1$ ($\gamma$ here just needs to be small, we verified the results for $\gamma=10^{-5}$ up to $\gamma=0.1$).
+This ensures that the training data is linearly separable in the first dimension of the data, but learning this first dimension requires having larger weight norm due to the scaling by $\gamma$.
 We leave the test data unchanged - our goal is to determine whether models trained on the modified training data can learn more than just the single identifying dimension.
-We then train logistic regression models to convergence (100 epochs) on both the reduced CIFAR-10 and CIFAR-100 tasks across a range of settings for weight decay, label smoothing, and Mixup.
+We then train logistic regression models on both the reduced CIFAR-10 and CIFAR-100 tasks across a range of settings for weight decay, label smoothing, and Mixup.
 We consider 20 uniformly spaced values in $[0,0.1]$ for the weight decay $\lambda$ parameter and in $[0,0.75]$ for the label smoothing $\alpha$ parameter, where the upper bound for the label smoothing parameter is obtained from the experiments of Müller et al.
@@
 (2018) and consider 20 uniformly spaced $\alpha$ values in $[0,8]$ (with $\alpha=0$ corresponding to ERM), which effectively covers the range of Mixup hyperparameter values used in practice.
-We train all models using Adam with the hyperparameters of $\beta_{1}=0.9,\beta_{2}=0.999$, a learning rate of $5*10^{-3}$, and a batch size of 500.
-After training, we report the mean test error along with one standard deviation error bounds over five training runs for each model setting in Figure 1.
-As can be seen from the results, both the label smoothing and Mixup models have high test error for all settings while weight decay achieves a significantly lower test error for all $\lambda>0$.
-The higher variance in the reported CIFAR-100 test errors can almost certainly be attributed to the fact that our binary version of CIFAR-100 consists of only 1000 training points and 200 points.
+Other hyperparameters are the same as before, except we use a learning rate of $5\times 10^{-3}$.
+The results across hyperparameter values are shown in Figure 2.
+Both the label smoothing and Mixup models have high test error for all settings while weight decay achieves a significantly lower test error for all $\lambda>0$.
+ERM also has high test error; in this case we differ from the setting of implicit bias results due to training with Adam and likely not training for the period required for convergence to the max-margin solution.
+Furthermore, these results are insensitive to introducing a small amount of weight decay to the label smoothing and Mixup models (i.e.
+$5\times 10^{-4}$ as in the previous subsection).
 (a) CIFAR-10
@@
 the ratio between the norm of the trained model weight vector in the first dimension and the remaining dimensions) for each of the trained models.
-4.2 Background Shift in Multi-Class Classification
-(a) Shaded MNIST
+4.3 Spurious Background Correlations
+(a) Colored MNIST
 (b) Model Performance
-Figure 3: Final model test errors over our hyperparameter sweep for the shaded MNIST dataset of Section 4.2, alongside a visualization of samples from the dataset.
-Although the experiments of Section 4.1 cleanly verify our theory, the spurious correlation introduced in the training data is effectively adversarial in nature.
-We now consider the case of a spurious correlation of greater practical interest: backgrounds that are correlated with the correct class label.
-In particular, we consider a very similar setup to the one used by Arjovsky et al.
-(2020) to motivate the influential invariant risk minimization framework: namely, we construct a shaded version of the MNIST dataset in which the background pixels for each class are replaced with different low intensity colors corresponding to the class labels.
+Figure 4: Final model test errors over our hyperparameter sweep for the colored MNIST dataset of Section 4.3, alongside a visualization of samples from the dataset.
+For our multi-class analogue to Section 4.2.1, we consider a similar setup to the one used by Arjovsky et al.
+(2020) to motivate the influential invariant risk minimization framework: namely, we construct a colored version of the MNIST dataset in which the background pixels for each class are replaced with different colors corresponding to the class labels.
 Unlike Arjovsky et al.
@@
 The only constraints we place on the background colors are that they are distinct across classes and that their intensities (i.e.
-values in RGB space) are small (it suffices to consider values bounded by 16); this makes it difficult to visually distinguish the background colors (hence "shaded MNIST") even though they are obviously different in RGB space.
-The color intensity constraint is not necessary for the failure of label smoothing and Mixup, but is necessary for weight decay to succeed since it ensures that the higher variance feature (the actual digit in the foreground) is generally more separable (due to larger pixel intensities).
+values in RGB space) are small (it suffices to consider values bounded by 16).
+The latter constraint is not necessary for the failure of label smoothing and Mixup, but is necessary for weight decay to succeed since it ensures that the higher variance feature (the actual digit in the foreground) is generally more separable (due to larger pixel intensities).
 Note that although there is no variance in the background color conditional on a class, there is significant variance in the actual pixels per class as the locations of the background pixels change across data points.
 For our model setup, we also follow Arjovsky et al.
-(2020) and consider a simple 2-layer feedforward neural network with ReLU activations and a hidden layer size of $2048$, as this is sufficient to achieve good performance on MNIST and is a tiny enough model that we can efficiently do the same hyperparameter sweep from Section 4.1.
-We follow the same training procedure as Section 4.1, training until the training loss roughly converges (only 20 epochs are necessary).
+(2020) and consider a simple 2-layer feedforward neural network with ReLU activations and a hidden layer size of $2048$, as this is sufficient to achieve good performance on MNIST and is a tiny enough model that we can efficiently do the same hyperparameter sweep from Section 4.2.1.
+Training details remain the same as in Section 4.2.1, except we only train for 20 epochs as this is sufficient for the training loss to roughly converge and greatly expedites the hyperparameter sweep.
 Test error results across the different hyperparameter settings for weight decay, label smoothing, and Mixup are shown in Figure 4.
@@
 5 Conclusion
-In this work, we have shown that label augmentation strategies such as label smoothing and Mixup exhibit a variance minimization effect that can lead to learning low variance, spurious correlations in the data even when there exists a better but higher variance feature.
-The main practical takeaway from this is that label smoothing and Mixup do not necessarily always improve the robustness of models, and their failure to do so may be indicative of certain identifying features in the training data (i.e.
-adversarially introduced patches or background pixels).
-Limitations.
-The main theoretical limitation of work is that our results for general models are only loss lower bounds and do not characterize optimization dynamics.
-Our experiments that such dynamics for feedforward networks would still reflect our observations, however, and it would be an interesting follow-up to study these dynamics using the techniques of Zou et al.
-(2023) and Chidambaram et al.
-(2023).
-On the empirical side, our main limitation is that we have demonstrated our theory using only small-scale experiments; a comprehensive empirical evaluation of label smoothing and Mixup in larger scale spurious correlations settings guided by our theory would be a natural avenue for future work.
-Broader Impacts.
-Although label augmentation are used to train and finetune large-scale models, our results concerning them in this work have mostly been theoretical.
+In this work, we have shown that label augmentation strategies such as label smoothing and Mixup exhibit a variance minimization effect (both in theory and in practice) that leads to lower variance intermediate representations and model outputs, which then correlate with better test performance on image classification benchmarks.
+A natural follow-up direction to our results is to investigate whether regularizers for encouraging lower variance features can be implemented directly to improve model performance, which would shed light on whether there is a causal relationship between this phenomenon and improved performance.
+Ethics Statement
+Although label smoothing and Mixup are used to train and fine-tune large-scale models, our results concerning them in this work have mostly been theoretical and explanatory.
 As a result, we do not anticipate any direct misuse of our results or any broader harmful impacts.
-Acknowledgements
-Rong Ge and Muthu Chidambaram are supported by NSF Award DMS-2031849 and CCF-1845171 (CAREER). Muthu would like to thank Annie Tang for thoughtful feedback during the early stages of this project.
+Reproducibility Statement
+All proofs of the results in this paper can be found in Appendix A. The supplementary material contains the code necessary to generate all figures, with instructions on how to run each experiment.
+Acknowledgments
+Rong Ge and Muthu Chidambaram were supported by NSF Award DMS-2031849 and CCF-1845171 (CAREER) during the completion of this work.
+Muthu would like to thank Annie Tang for thoughtful feedback during the early stages of this project.
 Appendix A Omitted Proofs
@@
 Lemma A.1.
-Let $\phi:\mathbb{R}^{d}\to\mathbb{R}$ be a twice-differentiable convex function satisfying $\gamma_{1}\mathrm{Id}\prec\nabla^{2}\phi\prec\gamma_{2}\mathrm{Id}$.
+Let $\phi:\mathbb{R}^{d}\to\mathbb{R}$ be a twice-differentiable convex function satisfying $\gamma_{1}\mathrm{I}_{d}\prec\nabla^{2}\phi\prec\gamma_{2}\mathrm{I}_{d}$.
 Then for any square integrable random variable $X$ on $\mathbb{R}^{d}$ it follows that:
@@
 Proof.
-Let us decompose the unique minimizer $w^{*}$ as $w^{*}=\alpha_{1}u^{*}+\alpha_{2}v^{*}$, where $u^{*}$ and $v^{*}$ are orthonormal and satisfy $u^{*}_{\mathcal{H}}=0,v^{*}_{\mathcal{L}}=0$ (i.e.
+Let us decompose the unique minimizer $w^{*}$ as $w^{*}=\alpha u^{*}+\beta v^{*}$, where $u^{*}$ and $v^{*}$ are orthonormal and satisfy $u^{*}_{\mathcal{H}}=0,v^{*}_{\mathcal{L}}=0$ (i.e.
 $u^{*}$ is a normalized version of the $\mathcal{L}$ components of $w^{*}$ and $v^{*}$ is the same but for the $\mathcal{H}$ components).
@@
 $(x,y)$ and decrease the loss of $w^{*}$ by replacing $v^{*}$ with $z^{*}$.
-Now suppose that $\alpha_{1}>\alpha_{2}$.
-Then we claim that replacing $\alpha_{1}$ and $\alpha_{2}$ with $\gamma=\sqrt{(\alpha_{1}^{2}+\alpha_{2}^{2})/2}$ yields a solution with lower loss than $w^{*}$.
-To see this, first observe that $2\gamma^{2}=\alpha_{1}^{2}+\alpha_{2}^{2}$, so that the norm of the modified solution is the same as $w^{*}$.
+Now suppose that $\alpha>\beta$. Then we claim that replacing $\alpha$ and $\beta$ with $\gamma=\sqrt{(\alpha^{2}+\beta^{2})/2}$ yields a solution with lower loss than $w^{*}$.
+To see this, first observe that $2\gamma^{2}=\alpha^{2}+\beta^{2}$, so that the norm of the modified solution is the same as $w^{*}$.
 This implies that the weight decay penalty term in $\ell_{\beta}$ is unchanged.
-Furthermore, we have that $\gamma\in(\alpha_{2},\alpha_{1})$, and that $\gamma-\alpha_{2}>\alpha_{1}-\gamma$.
-This follows from the fact that $\gamma\geq(\alpha_{1}+\alpha_{2})/2$ by construction.
+Furthermore, we have that $\gamma\in(\beta,\alpha)$, and that $\gamma-\beta>\alpha-\gamma$.
+This follows from the fact that $\gamma\geq(\alpha+\beta)/2$ by construction.
 This, combined with the fact that $y\left\langle v^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$, then implies for $\pi$-a.e.
 $(x,y)$ that:
-$y\left\langle\gamma u^{*}+\gamma v^{*},x\right\rangle>y\left\langle\alpha_{1}u^{*}+\alpha_{2}v^{*},x\right\rangle.$ (A.2)
-Which contradicts the optimality of $w^{*}$. Therefore, we must have $\alpha_{1}\leq\alpha_{2}$, which gives the desired result. ∎
+$y\left\langle\gamma u^{*}+\gamma v^{*},x\right\rangle>y\left\langle\alpha u^{*}+\beta v^{*},x\right\rangle.$ (A.2)
+Which contradicts the minimality of $w^{*}$. Therefore, we must have $\alpha\leq\beta$, which gives the desired result. ∎
 See 3.4
 Proof.
-The result follows from the same decomposition approach of Theorem 3.4 and the definition of the max-margin solution (Definition 2.1).
-Namely, we can move weight from the max-margin direction in $\mathcal{L}$ to the max-margin direction in $\mathcal{H}$ to decrease the norm while satisfying the separation constraint.
+We can apply the same decomposition as in the proof of Theorem 3.4; namely, $w^{*}=\alpha u^{*}+\beta v^{*}$ where $w^{*}$ is the max-margin solution and $u^{*}$ and $v^{*}$ are as before.
+Suppose again that $\alpha>\beta$ and let $\gamma=\sqrt{(\alpha^{2}+\beta^{2})/2}$.
+We claim that there exists $\epsilon\in(0,\gamma)$ such that $w=(\gamma-\epsilon)u^{*}+\gamma v^{*}$ satisfies $y\left\langle w,x,\geq\right\rangle 1$ for $\pi$-a.e.
+$(x,y)$, which would contradict $w^{*}$ being the max-margin solution since $\norm{w}^{2}<\norm{w^{*}}^{2}$.
+From the exact same logic as in the proof of Theorem 3.4, we have:
+$y\left\langle\gamma u^{*}+\gamma v^{*},x\right\rangle-y\left\langle w^{*},x\right\rangle>0$ (A.3)
+for $\pi$-a.e.
+$(x,y)$.
+Since $\mathrm{supp}(\pi)$ is compact, (A.3) attains a minimum $\kappa>0$ over $\mathrm{supp}(\pi)$.
+Similarly, $y\left\langle u^{*},x\right\rangle$ attains a finite maximum.
+By choosing $\epsilon$ such that $y\left\langle\epsilon u^{*},x\right\rangle<\kappa$, we obtain the desired contradiction.
 ∎
@@
 Once again, since we can restrict our attention to $g(X)\in[\gamma,1-\gamma]$ for some $\gamma$ (as the loss goes to infinity if $g(X)=1$ on a set of positive $\pi_{X}$-measure with $\alpha>0$), it is easy to verify that (A.13) is strongly convex in $g(X)$.
-The desired result then follows by applying Lemma A.1 with the regular conditional distribution $g(X)\mid Y=y$ for each term in (A.13).
+The desired result then follows by applying Lemma A.1 with the regular conditional distribution $\pi_{X\mid Y=y}$ for each term in (A.14).
 ∎
@@
 ∎
+Appendix B Comparison to Settings of Prior Work
+Here we review the settings of Chidambaram et al. (2023) and Zou et al. (2023) in greater detail to provide a more precise comparison to the setting in our work.
+Setting of Chidambaram et al.
+(2023).
+The authors consider a multi-view data model inspired by Allen-Zhu & Li (2021); namely, their data distribution $\pi$ is a multi-class distribution supported on $\mathbb{R}^{Pd}\times[k]$ where $P$ corresponds to the number of patches in each data point.
+Essentially, for $(x,y)\sim\pi$, we view $x$ as $x=\{x^{(1)},x^{(2)},...,x^{(P)}\}$ with each $x^{(i)}\in\mathbb{R}^{d}$.
+The purpose of this partitioning is so that features related to the target class can appear in some patches and noise can appear in the remaining patches.
+In particular, the authors consider two target features per class ($v_{y,1}$ and $v_{y,2}$) that appear in a constant number of signal patches, while all other patches in a data point $x$ correspond to low magnitude feature noise, i.e.
+these patches consist of some linear combination of features $v_{s,j}$ for $s\neq y$.
+Furthermore, each signal patch has only a single feature (either $v_{y,1}$ or $v_{y,2}$) with a random weight $\beta$ such that for any signal patch $x^{(p)}=\beta v_{y,1}$ there is another patch $x^{(q)}=(C-\beta)v_{y,2}$ for a fixed parameter $C$.
+The authors emphasize that a model that has the same correlation with both $v_{y,1}$ and $v_{y,2}$ will achieve lower variance, as it will have a constant total correlation and be insensitive to the variation in $\beta$.
+For this type of data distribution, the authors consider the training dynamics of a two-layer convolutional neural network with smoothed ReLU activations and non-trainable second layer weights (not an issue in this case, since second layer weights can be absorbed into the first layer).
+They analyze training using the empirical cross-entropy, as well as the Mixup cross-entropy for the specific case of a mixing distribution $\mathcal{D}_{\lambda}$ that is just a point mass on $1/2$ (Midpoint Mixup).
+The authors prove that, running gradient descent on the empirical cross-entropy leads (with high probability) to a model that only learns one feature for almost all classes, while doing the same for Midpoint Mixup yields a model that learns both features per class.
+The results are asymptotic; the authors consider all hyperparameters in their setup to be sufficiently large (even the number of classes $k$) or small.
+Setting of Zou et al.
+(2023).
+Similar to Chidambaram et al.
+(2023), Zou et al.
+(2023) also works in a setting motivated by Allen-Zhu & Li (2021) and consider data consisting of patches.
+However, they focus on binary classification, and their data distribution $\pi$ is supported on $\mathbb{R}^{Pd}\times\{1,2\}$.
+The data $x$ has exactly one (randomly selected) patch $x^{(i)}$ that contains a target feature; Zou et al.
+(2023) delineate one type of target feature $v$ as common and another type $v^{\prime}$ as rare.
+There are also up to $b$ other patches in $x$ that consist of common features from other classes (i.e.
+feature noise, like in Chidambaram et al.
+(2023)).
+Lastly, different from Chidambaram et al.
+(2023), Zou et al.
+(2023) consider the leftover patches to consist of i.i.d.
+Gaussian noise.
+Zou et al.
+(2023) also consider the training dynamics of a two-layer convolutional neural network with frozen second layer weights on this type of data distribution, but use a squared activation instead of a smoothed ReLU.
+Unlike the results of Chidambaram et al.
+(2023) which are stated entirely in terms of whether the features $v_{y,1}$ and $v_{y,2}$ are learned in a certain sense, Zou et al.
+(2023) directly prove a lower bound on the test error of models trained using gradient descent on the empirical cross-entropy while also showing that the test error of models trained on the empirical Mixup cross-entropy is vanishing small at some time step during model training.
+Essentially, this is due to the non-Mixup models failing to learn the rare feature $v^{\prime}$ for both classes.
+Their results apply to Mixup with a mixing distribution $\mathcal{D}_{\lambda}$ that is any point mass on $(0.5,1)$.
+Differences in our setting.
+Both Chidambaram et al.
+(2023) and Zou et al.
+(2023) prove results concerning gradient descent dynamics, whereas our results directly consider the minimizers of the population losses associated with weight decay, label smoothing, and Mixup.
+Here our choice to work with the population losses is not substantially different from Chidambaram et al.
+(2023) and Zou et al.
+(2023), since although they work with the empirical losses they are in an asymptotic setting and large swaths of the proofs in both papers rely on concentration of measure arguments.
+However, we do differ substantially in that our main results apply to linear models – this makes our results less practical but technically much simpler than the results of Chidambaram et al.
+(2023) and Zou et al.
+(2023), with the added benefit that we can also handle any symmetric mixing distribution $\mathcal{D}_{\lambda}$.
+Furthermore, since our main results work in this linear setting, we also adopt a much simpler data distribution setup.
+We do not consider our input data $x$ as partitioned into patches, instead we merely designate some subset of the input dimensions as “low variance” and the complementary subset as “high variance”.
+In this sense, we do not have explicit feature vectors associated with each class $y$ like the $v_{y,1},v_{y,2}$ of Chidambaram et al.
+(2023) or the $v,v^{\prime}$ of Zou et al.
+(2023).
+Appendix C Additional Experiments
+C.1 Visualization of Decision Boundaries
+(a) Canonical Hyperparameters
+(b) Scaled Hyperparameters (1)
+(c) Scaled Hyperparameters (2)
+Figure 5: Visualization of weight decay, label smoothing, and Mixup decision boundaries.
+Figure (a) considers canonical hyperparameter choices, and Figures (b) and (c) illustrate the effects of scaling these choices.
+To provide some intuition for Theorems 3.3 to 3.7, we visualize the decision boundaries of trained logistic regression models on 2-D data.
+In particular, denoting the classes as usual by $y\in\{-1,+1\}$, we consider a simple data distribution in which the first coordinate is distributed uniformly on $[y,10y]$ and the second coordinate is fixed to be $0.1y$.
+This data distribution is linearly separable in each coordinate; however, the second coordinate is fixed and thus has no conditional variance.
+Consequently, Theorems 3.6 and 3.7 predict that minimizing the population label smoothing and Mixup losses on this data should lead to learning a model whose decision boundary is aligned with the $x$-axis (i.e.
+we only use the second coordinate to determine which class we predict).
+To verify this, we visualize the decision boundaries of logistic regression models trained on $500$ points sampled from this distribution using multiple different settings of weight decay, label smoothing, and Mixup in Figure 5.
+We train for $500$ epochs using full batch SGD with a learning rate of $10^{-2}$, although our results were not sensitive to these choices.
+We use a mixing distribution of $\mathrm{Beta}(\alpha,\alpha)$ for Mixup (as is standard), and consider the canonical hyperparameter choices of $5\times 10^{-4}$ for weight decay, $0.1$ for label smoothing, and $\mathrm{Beta}(1,1)$ for Mixup (Zhang et al., 2018; Müller et al., 2020) in Figure 5 (a) and also check the effect of scaling these hyperparameters in Figure 5 (b) and (c).
+As can be seen from the results, the label smoothing and Mixup decision boundaries are much closer to being aligned with the $x$-axis than the weight decay boundary, with the alignment getting stronger with more extreme choices of the hyperparameters.
+The fact that the boundaries are not exactly aligned with the $x$-axis can be attributed to the fact that we are not exactly minimizing the population loss, since we are in the finite sample, fixed training horizon setting.
+That the label smoothing boundary is more aligned to the $x$-axis than the Mixup boundary is also to be expected, since the Mixup loss introduces randomness in the form of the mixing distribution which makes it more difficult to minimize.
+Lastly, the fact that the boundaries for both label smoothing and Mixup become more aligned with the $x$-axis at more extreme hyperparameter values can be intuitively explained by the fact that the loss incurred by predicting a probability close to 1 for either class increases as we scale the hyperparameters, i.e.
+we suffer more loss from relying on the high variance first coordinate.
+Remark C.1.
+We should point out that, while our theoretical and empirical results in Figure 5 show a clear difference in the types of solutions learned when training linear classifiers using label smoothing, Mixup, and weight decay, it is not always the case that the augmented losses lead to different solutions than ERM (with possibly weight decay).
+Indeed, Chidambaram et al.
+(2021) and Oh & Yun (2023) both showed that for different settings of Gaussian data, ERM and Mixup can lead to learning the same solution.
+However, the settings of these prior works don’t fall within our scope as we consider distributions with compact support, which ends up being an important property for proving our results.
+C.2 Direct Verification of Theory
+Here we directly analyze the training of logistic regression models on a synthetic data distribution that exactly follows the assumptions of Definition 3.1; the following definition generalizes the 2-D distribution we visualized in Section C.1.
+Definition C.2 (Synthetic Data).
+We define a distribution $\pi_{\gamma}$ parameterized by $\gamma$ (where $10<\gamma<$ 1) on $\mathbb{R}^{d}\times\{-1,1\}$ with $\pi_{\gamma,Y}(1)=1/2$ and $x\sim\pi_{\gamma,X\mid Y=y}$ satisfying:
+  1. 1.
+    First $\lfloor d/2\rfloor$ dimensions are constant but small. We have $x_{i}=\gamma y$ for $i\leq\lfloor d/2\rfloor$.
+  2. 2.
+    Last $d-\lfloor d/2\rfloor$ dimensions are high variance. We have $x_{i}\sim\mathrm{Uniform}([y,100y])$ for $i>\lfloor d/2\rfloor$.
+In other words, we consider data where (conditional on the label) the first half of the dimensions are fixed (corresponding to $\mathcal{L}$) and the second half are i.i.d.
+high variance uniform (corresponding to $\mathcal{H}$).
+We sample $n=5000$ data points according to Definition C.2 with $d=10$ and $\gamma=0.1$ and train logistic regression models across a range of settings for weight decay, label smoothing, and Mixup.
+The choice of $\gamma$ here, as well as the range of values for the dimensions in $\mathcal{H}$, is relatively arbitrary; we verified our empirical results hold for different scales of $\gamma$ and ranges for $\mathcal{H}$.
+The empirical results also do not depend on the fact that the $\mathcal{L}$ coordinates are zero variance – we checked that they still hold when adding a small magnitude uniform noise to the $\mathcal{L}$ coordinates.
+For model training, we consider 20 uniformly spaced values in $[0,0.1]$ for the weight decay $\lambda$ parameter and in $[0,0.75]$ for the label smoothing $\alpha$ parameter, where the upper bound for the label smoothing parameter space is obtained from the experiments of Müller et al.
+(2020).
+For Mixup, we fix the mixing distribution to be the canonical choice of $\mathrm{Beta}(\alpha,\alpha)$ introduced by Zhang et al.
+(2018) and consider 20 uniformly spaced $\alpha$ values in $[0,8]$ (with $\alpha=0$ corresponding to ERM), which effectively covers the range of Mixup hyperparameter values used in practice.
+We train all models for 100 epochs using AdamW with the standard hyperparameters of $\beta_{1}=0.9,\beta_{2}=0.999$, a learning rate of $5\times 10^{-3}$, and a batch size of 500.
+At the end of training, we compute $\norm{w_{\mathcal{H}}}$ (i.e.
+the norm of the weight vector in the last 5 dimensions) for each trained model.
+For each model setting, we report the mean and standard deviation of $\norm{w_{\mathcal{H}}}$ over 5 training runs in Figure 6.
+(a) Weight Decay
+(b) Label Smoothing
+(c) Mixup
+Figure 6: Final weight norm of logistic regression model in the high variance dimensions ($\norm{w_{\mathcal{H}}}$) for various hyperparameter settings on the synthetic data distribution of Section C.2.
+As can be seen from the results, the weight decay models always have non-trivial values for $\norm{w_{\mathcal{H}}}$ (even for large values of $\lambda$), whereas the label smoothing and Mixup models very quickly converge to a $\norm{w_{\mathcal{H}}}$ of effectively zero as their respective hyperparameters move away from the ERM regime ($\alpha=0$ in both cases).
+This matches the behavior predicted by Theorems 3.3 to 3.7.
+C.3 Full ResNet Results
+Here we collect final test error/variance results for the plots shown in Section 4.1 and also provide analogous plots and results for ResNet-50 and ResNet-101 models.
+In all of the following, we abbreviate weight decay as “WD” and label smoothing as “LS”.
+C.3.1 ResNet-18 Final Results
+Tables 1 and 2 show the end-of-training test error, activation variance, and output variance results for the experiments in Figure 1.
+Method Test Error Activation Variance Output Variance
+ERM + WD $8.40\pm 0.58$ $198\pm 5$ $0.133\pm 0.008$
+ERM + WD + LS $7.96\pm 0.30$ $33.0\pm 5.9$ $0.099\pm 0.003$
+ERM + WD + Mixup $\mathbf{5.90}\pm 0.21$ $18.4\pm 0.4$ $0.059\pm 0.002$
+Table 1: Final results (mean test error/variance and one standard deviation over 5 runs) for ResNet-18 experiments on CIFAR-10.
+Method Test Error Activation Variance Output Variance
+ERM + WD $31.59\pm 0.36$ $673\pm 23$ $0.406\pm 0.007$
+ERM + WD + LS $28.55\pm 4.7$ $54.7\pm 5.6$ $0.188\pm 0.027$
+ERM + WD + Mixup $\mathbf{26.83}\pm 0.44$ $82.2\pm 1.2$ $0.143\pm 0.004$
+Table 2: Final results (mean test error/variance and one standard deviation over 5 runs) for ResNet-18 experiments on CIFAR-100.
+C.3.2 ResNet-50 All Results
+(a) CIFAR-10 Test Error
+(b) CIFAR-10 Activation Variance
+(c) CIFAR-10 Output Variance
+(d) CIFAR-100 Test Error
+(e) CIFAR-100 Activation Variance
+(f) CIFAR-100 Output Variance
+Figure 7: ResNet-50 final test errors, penultimate layer activation variances, and output probability variances on CIFAR-10 and CIFAR-100.
+We train ResNet-50 models under the same hyperparameters as in Section 4.1, except for a batch size of 512 due to memory constraints.
+ResNet-50 results analogous to those shown in Figure 1 are shown in Figure 7.
+Similarly, final error and variance results are shown in Tables 3 and 4.
+We observe that while the same trends hold as in the case of ResNet-18, there is significantly more variance in the computed activation variances for Mixup on CIFAR-10.
+In this particular case, we found that there were still significant oscillations in activation variances even towards the end of the training horizon.
+This may in part be attributable to reduced batch size, but we did not investigate this further.
+Method Test Error Activation Variance Output Variance
+ERM + WD $7.36\pm 0.13$ $551\pm 711$ $0.119\pm 0.002$
+ERM + WD + LS $7.24\pm 0.26$ $6.35\pm 3.47$ $0.091\pm 0.004$
+ERM + WD + Mixup $\mathbf{5.41}\pm 0.37$ $95\pm 115$ $0.054\pm 0.003$
+Table 3: Final results (mean test error/variance and one standard deviation over 5 runs) for ResNet-50 experiments on CIFAR-10.
+Method Test Error Activation Variance Output Variance
+ERM + WD $29.41\pm 0.94$ $310\pm 14$ $0.383\pm 0.012$
+ERM + WD + LS $28.58\pm 0.28$ $13.14\pm 0.20$ $0.257\pm 0.001$
+ERM + WD + Mixup $\mathbf{24.94}\pm 1.63$ $22.95\pm 1.52$ $0.161\pm 0.007$
+Table 4: Final results (mean test error/variance and one standard deviation over 5 runs) for ResNet-50 experiments on CIFAR-100.
+C.3.3 ResNet-101 All Results
+(a) CIFAR-10 Test Error
+(b) CIFAR-10 Activation Variance
+(c) CIFAR-10 Output Variance
+(d) CIFAR-100 Test Error
+(e) CIFAR-100 Activation Variance
+(f) CIFAR-100 Output Variance
+Figure 8: ResNet-101 final test errors, penultimate layer activation variances, and output probability variances on CIFAR-10 and CIFAR-100.
+We also train ResNet-101 models under the same hyperparameters as in Section 4.1, except once again for a batch size of 512 due to memory constraints.
+ResNet-101 results analogous to those shown in Figure 1 are shown in Figure 8.
+Similarly, final error and variance results are shown in Tables 5 and 6.
+Here we see that the variance oscillation behavior that we mentioned in Appendix C.3.2 is even more pronounced for the CIFAR-10 results, suggesting that this behavior is amplified for larger models.
+Once again, it is not clear what properties of CIFAR-10 lead to highly oscillatory activation variances for some initializations, but we again suspect that reduced batch size in training at least plays some role.
+The CIFAR-100 results remain consistent, although there is still some oscillatory behavior for the Mixup results.
+Method Test Error Activation Variance Output Variance
+ERM + WD $6.87\pm 0.17$ $491\pm 235$ $0.110\pm 0.002$
+ERM + WD + LS $6.90\pm 0.26$ $2280\pm 2292$ $0.088\pm 0.003$
+ERM + WD + Mixup $\mathbf{5.23}\pm 0.18$ $1701\pm 1586$ $0.054\pm 0.002$
+Table 5: Final results (mean test error/variance and one standard deviation over 5 runs) for ResNet-101 experiments on CIFAR-10.
+Method Test Error Activation Variance Output Variance
+ERM + WD $28.69\pm 0.73$ $283\pm 18$ $0.373\pm 0.008$
+ERM + WD + LS $27.74\pm 0.55$ $12.01\pm 0.23$ $0.269\pm 0.005$
+ERM + WD + Mixup $\mathbf{23.55}\pm 0.98$ $115\pm 136$ $0.157\pm 0.01$
+Table 6: Final results (mean test error/variance and one standard deviation over 5 runs) for ResNet-101 experiments on CIFAR-100.
+C.4 Output Probability Variance Analysis
+A natural explanation for why output probability variance decreases for label smoothing and Mixup in Figures 1, 7, and 8 is that label smoothing and Mixup improve test error and consequently have less variability in their outputs due to fewer mistakes.
+Firstly, looking carefully at the label smoothing results in the previous subsections shows this cannot be the full cause, as in Table 5 label smoothing leads to worse test error than the baseline of ERM + WD but still leads to lower output probability variance.
+In Tables 7 to 9, we compute the average output probability variance for each class but consider only the probability associated with the target class (i.e.
+for all points with label $y$, we compute the variance of the predicted probabilities corresponding to $y$).
+In all cases, the target output variance alone leaves a significant fraction of the overall output variance unexplained.
+Method Target Output Variance (CIFAR-10) Target Output Variance (CIFAR-100)
+ERM + WD $0.066\pm 0.004$ $0.172\pm 0.002$
+ERM + WD + LS $0.050\pm 0.002$ $0.119\pm 0.002$
+ERM + WD + Mixup $0.029\pm 0.002$ $0.087\pm 0.002$
+Table 7: Target output probability variance for ResNet-18.
+Method Target Output Variance (CIFAR-10) Target Output Variance (CIFAR-100)
+ERM + WD $0.059\pm 0.001$ $0.166\pm 0.004$
+ERM + WD + LS $0.046\pm 0.002$ $0.131\pm 0.001$
+ERM + WD + Mixup $0.027\pm 0.002$ $0.093\pm 0.004$
+Table 8: Target output probability variance for ResNet-50.
+Method Target Output Variance (CIFAR-10) Target Output Variance (CIFAR-100)
+ERM + WD $0.054\pm 0.001$ $0.162\pm 0.002$
+ERM + WD + LS $0.044\pm 0.001$ $0.131\pm 0.002$
+ERM + WD + Mixup $0.026\pm 0.002$ $0.088\pm 0.004$
+Table 9: Target output probability variance for ResNet-101.